### v1

Abstract
Differentially private distributed mean estimation (DP-DME) is a fundamental building block in privacy-preserving federated learning, where a central server estimates the mean of $d$-dimensional vectors held by $n$ users while ensuring $(\epsilon,\delta)$-DP. Local differential privacy (LDP) and distributed DP with secure aggregation (SecAgg) are the most common notions of DP used in DP-DME settings with an untrusted server. LDP provides strong resilience to dropouts, colluding users, and malicious server attacks, but suffers from poor utility. In contrast, SecAgg-based DP-DME achieves an $O(n)$ utility gain over LDP in DME, but requires increased communication and computation overheads and complex multi-round protocols to handle dropouts and malicious attacks. In this work, we propose CorDP-DME, a novel DP-DME mechanism that spans the gap between DME with LDP and distributed DP, offering a favorable balance between utility and resilience to dropout and collusion. CorDP-DME is based on correlated Gaussian noise, ensuring DP without the perfect conditional privacy guarantees of SecAgg-based approaches. We provide an information-theoretic analysis of CorDP-DME, and derive theoretical guarantees for utility under any given privacy parameters and dropout/colluding user thresholds. Our results demonstrate that (anti) correlated Gaussian DP mechanisms can significantly improve utility in mean estimation tasks compared to LDP – even in adversarial settings – while maintaining better resilience to dropouts and attacks compared to distributed DP.
[…]
1 Introduction
Distributed mean estimation (DME) is a fundamental building block in a number of applications ranging from federated learning [1, 2], distributed stochastic gradient descent [3, 4, 5, 6] to distributed sensor network computations [7]. Differentially private DME (DP-DME) refers to the setting where a central server aims to estimate the mean of $d$ dimensional vectors held by $n$ distributed users while ensuring differential privacy (DP) [8, 9] of the users’ vectors. The utility of DP-DME is measured by the mean squared error (MSE) between the estimate at the server and the true mean. DP-DME has been studied under multiple notions of DP: (i) Central DP (CDP) [10], where a trusted server collects user vectors, computes the mean, and adds noise to ensure DP; (ii) Local DP (LDP), where each user independently perturbs their vector before sending to a potentially malicious server, and (iii) Distributed DP [11], where users perturb vectors locally and use secure aggregation (SecAgg) [12, 13] to ensure the server only views the sum of perturbed vectors [14, 15, 16, 17].
LDP and distributed DP using SecAgg are the most common approaches for DME that avoid the requirement of a trusted server – see Table 1 for a comparison. DME with LDP requires nearly no coordination among users and, consequently, maintains privacy even if an arbitrary number of users collude with the server. Additionally, it provides a graceful degradation of utility with user dropouts – that is, the MSE increases gradually as the number of users who drop out increases. However, the privacy resilience of LDP-based DME comes at a large utility cost: relative to CDP, LDP-based DME increases MSE by a factor of $O(n)$ [18].
Distributed DP via SecAgg, on the other hand, requires perfect coordination across users and a complex multi-party cryptographic protocol performed over multiple rounds. This approach has two key features: (i) it ensures that the server learns only the sum of the user’s vectors (and nothing else) and (ii) it allows for an MSE that is achieved by CDP, and hence an $O(n)$ factor reduction in the MSE as compared to LDP. In effect, because of (i), each user only needs to add enough noise to ensure privacy of the summed vectors, which enables the server’s input – as well as the resulting MSE – to be statistically identical to the CDP case. However, the improvement in MSE comes at a coordination cost. SecAgg requires perfectly (anti) correlated random noise from a finite field to mask the users’ vectors to ensure that the server can only access the sum of the users’ vectors. This makes the method less resilient to dropouts in terms of utility and vulnerable to malicious server attacks that involve fake client registrations and strategic connection drops [19]. Notably, the protocol can be made resilient to dropouts and collusions up to fixed thresholds on the number of dropouts and colluding users. However, unlike LDP, this resilience requires multiple rounds of communication and increased shared randomness among users. Moreover, when the collusion and dropout thresholds are exceeded, DP-DME with SecAgg results in sudden drops in privacy and utility, respectively.
In this work, we develop novel DP-DME mechanisms that simultaneously retain the key advantages of both LDP and SecAgg. Specifically, we target techniques that: (i) like LDP, can be performed with one round without requiring complex cryptographic approaches to handle dropouts and collusions, (ii) like SecAgg, provide MSE comparable to CDP, (iii) are resilient like LDP, in that they yield graceful degradation of utility with increasing number of collusions and dropouts irrespective of the thresholds, (iv) maintains the required level of privacy up to the colluding user threshold as in both LDP and distributed DP with SecAgg, and graceful decay of privacy beyond the threshold. State-of-the-art distributed DP methods for DME use cryptographic protocols such as SecAgg that operate over a finite field and ensure perfect privacy of the inputs and perfect reconstruction of the sum of the inputs. However, these methods ultimately provide DP guarantees to users’ real-valued vectors and approximate reconstructions of the mean. We introduce an alternative DP-DME setting that is end-to-end differentially private.
Our starting point (see Sec. 3) is a novel interpretation of one of the core ideas of SecAgg: utilizing (anti) correlated noise to reduce the overall MSE without compromising on privacy. Unlike SecAgg, which requires correlated noise over a finite field, we consider adding correlated Gaussian noise directly to real-valued vectors held by users.
Figure 1: MSE with different numbers of responding users for $n=100$, $\epsilon=2$, $\delta=10^{-5}$. CorDP-DME spans the gap between DME with LDP and CDP.
We couple this starting point with a DP framework for quantifying the impact of arbitrary noise correlations on privacy, particularly in face of user dropouts and collusion. We apply this framework to develop a simple, one-round DP-DME protocol called CorDP-DME. After an initial offline phase to establish shared randomness, CorDP-DME uses an optimized correlated Gaussian mechanism among users and provides the utility of CDP without resorting to SecAgg (See Table 1). In the presence of dropouts and collusions, CorDP-DME achieves significantly reduced MSE relative to LDP yet retains its flexibility, resilience, and simplicity as compared to SecAgg. Notably, CorDP-DME spans the gap between the two extremes of LDP and distributed DP with SecAgg (Fig. 1).
1.1 Our Contributions
  • We introduce a DP-DME model that allows users to execute privacy mechanisms that are arbitrarily correlated with each other. The proposed model assumes a malicious server that attempts to compromise users’ privacy, and considers dropouts and colluding users. The generalized notion of DP for DME introduced in this work covers both LDP and distributed DP with SecAgg as special cases (Section 2).
  • We provide a novel geometric interpretation of the use of (anti) correlated privacy mechanisms in DP-DME, which demonstrates how correlated mechanisms can be used to achieve utility gains over independent noise mechanisms used by LDP (Section 3).
  • We perform an information-theoretic analysis on the correlated Gaussian mechanism for DP-DME to determine the optimum noise parameters and the decoder at the server that minimizes the MSE for any given privacy parameters and dropout/colluding user thresholds (Section 5).
  • We propose CorDP-DME, a single-round DP-DME mechanism that spans the gap between DME with LDP and distributed DP with SecAgg in terms of privacy-utility trade-offs and resilience to dropouts and attacks (Sections 4 and 5).
LDP [20] Distributed DP w/ SecAgg [12, 16] CorDP-DME
Communication Single round Multiple rounds Single round
Dropouts Gradual rise in MSE with dropouts Gradual rise in MSE with dropouts up to the threshold, MSE surge afterwards Gradual rise in MSE with dropouts
Collusion $(\epsilon,\delta)$-DP with any number of colluding users $(\epsilon,\delta)$-DP up to $c<n/3$ colluding users, sudden drop in privacy afterwards $(\epsilon,\delta)$-DP up to any $c$ colluding users, graceful privacy decay afterwards
MSE:no dropouts $O\left(\frac{d}{n\min\{\epsilon,\epsilon^{2}\}}\right)$ $O\left(\frac{d}{n^{2}\min\{\epsilon,\epsilon^{2}\}}\right)$ $O\left(\frac{d}{n^{2}\min\{\epsilon,\epsilon^{2}\}}\right)$
[…]
Computation User: $O(d)$, Server: $O(d)$ User: $O(n^{2}+dn)$, Server: $O(dn^{2})$ User: $O(dn)$, Server: $O(d)$
Table 1: Comparison of CorDP-DME with existing approaches for DP-DME.
1.2 Related Work
LDP-DME:
As the user has the complete control over the perturbations made to their private vectors, LDP-DME assumes the strongest threat model in DP-DME. Optimality results on LDP-DME have been provided in [21] for $\epsilon\leq 1$ (asymptotic) and in [20, 22] for any $\epsilon$, based on the lower bounds provided in [18]. A number of order optimal LDP-DME algorithms and fundamental results on communication constraints have been provided in [23, 24, 25, 26, 27, 25, 28, 29, 30]. CorDP-DME fundamentally differs from LDP by allowing for the privacy mechanisms among different users to be arbitrarily correlated. In fact, the system model introduced in this work is a generalization of additive DME mechanisms with LDP.
Correlated noise in DP:
Our work is closely related to [31], where the server has access to the perturbed vectors of all users in the DME setting, which are correlated. The DP-DME mechanism introduced in [31] uses a noise generation protocol that is based on SecAgg [12], and achieves the same level of utility as CDP-DME for certain values of $\epsilon$ and $\delta$. Our work deviates from this setting in two main aspects: (i) Dropouts and collusion threshold: CorDP-DME designs the correlated noise distributions for any given thresholds on dropouts and colluding users without bounds. In [31], dropouts are not considered and the collusion threshold is bounded ($<n/3$); (ii) Difference in the problem formulation: The noise generation protocol in [31] uses a fixed correlation coefficient of $-\frac{1}{n}$ among the privacy mechanisms used by different users, that specifies achievable pairs of $(\epsilon,\delta)$ for a target MSE. In this work, we determine the optimum noise distribution (including the correlation coefficient) for any given $\epsilon,\delta,n,c,u$ to minimize the MSE for the required level of privacy.
The advantage of using correlated privacy mechanisms in decentralized learning settings with graph-based approaches is studied in [32, 33]. In these references, the users are represented as nodes in a graph with each edge representing a communication link. The neighboring users (that share an edge) add perfectly (anti) correlated noise to their updates in such a way that they cancel out when averaged. An independent noise component is also added to ensure the privacy of the aggregated updates. The problem setting in [32] differs from ours primarily by not considering the effects of dropouts after noise initialization. In reference [33], dropouts are handled in three additional rounds of user-user communications, unlike CorDP-DME which requires no additional rounds. The line of works related to differentially private follow-the-regularized leader (DP-FTRL) [34, 35, 36, 37] investigate the role of correlated noise across iterations in private learning algorithms.
2 Problem Formulation
We consider a distributed mean estimation (DME) setting with $n$ users, each holding an independent $d$-dimensional vector, and a central server that estimates the mean of these vectors while ensuring $(\epsilon,\delta)$-DP. We assume $n$ point-to-point links, one link from each user to the server in the underlying communication network. Let $[1:n]\coloneq\{1,\dotsc,n\}$ denote the set of all users. Each user $i\in[1:n]$ generates a vector $\mathbf{x}_{i}\in\mathbb{B}^{d}$, where $\mathbb{B}^{d}\subset\mathbb{R}^{d}$ is the unit ball, and sends a distorted version
$M(\mathbf{x}_{i})\coloneq\mathbf{x}_{i}+\mathbf{Z}_{i}.$ (1)
to the server, where $\mathbf{Z}_{i}\in\mathbb{R}^{d}$ are (not necessarily independent) random noise variables. We assume a potentially malicious server and allow for up to a given number of colluding users and dropouts. Let $\mathcal{C}\subset[1:n]$, $|\mathcal{C}|\leq c$ denote the set of users colluding with the server, i.e., the server has access to both $\mathbf{x}_{i}$ and $M(\mathbf{x}_{i})$ for $i\in\mathcal{C}$. Moreover, let $\mathcal{F}\subset[1:n]$, $|\mathcal{F}|\geq t$ be the set of responding users at a given time. We assume $c<t$. The set of non-colluding and responding users is denoted by $\mathcal{U}=\mathcal{F}\setminus\mathcal{C}$.
The server decodes the mean of the vectors of the responding users in $\mathcal{U}$ using a linear function $d_{\mathcal{U}}:\mathbb{R}^{d|\mathcal{U}|}\rightarrow\mathbb{R}^{d}$, defined as:
$\hat{\mathbf{S}}_{\mathcal{U}}$ :-$\coloneq d_{\mathcal{U}}([M(\mathbf{x}_{j_{1}}),\dotsc,M(\mathbf{x}_{j_{|\mathcal{U}|}})])=\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\alpha_{i}M(\mathbf{x}_{i}),:$ (2)
where $\mathcal{U}=\{j_{1},\dotsc,j_{|\mathcal{U}|}\}$, $j_{1}<j_{2}<\dotsc<j_{|\mathcal{U}|}$, and $\alpha_{i}\in\mathbb{R}$, $i\in\mathcal{U}$ are constants (for a given $\mathcal{F}$ and $\mathcal{C}$) that define the decoding function. Ideally, $\hat{\mathbf{S}}_{\mathcal{U}}$ is approximately the mean of the vectors held by responding users.
We consider a secure DME protocol that consists of an offline (data independent) phase and an online (data dependent) phase. In the offline phase, the users exchange information amongst themselves to ensure that user $i$ has random variable $\mathbf{Z}_{i}$ where $(\mathbf{Z}_{1},\mathbf{Z}_{2},\ldots,\mathbf{Z}_{n})$ have a joint distribution $\pi$. In the online phase, user $i$ sends $M(\mathbf{x}_{i})$ to the server as per (1), and the server outputs $\hat{\mathbf{S}}_{\mathcal{U}}$. Let $\mathcal{I}_{j}$ denote all the information used to generate $\mathbf{Z}_{j}$ for $j\in\mathcal{C}$ in the offline phase. In the DME protocol, after both the offline and online phases, the server has access to $\{\mathcal{I}_{j},\mathbf{x}_{j},\mathbf{Z}_{j}\}_{j\in\mathcal{C}}$, and $\{M(\mathbf{x}_{j})\}_{j=1}^{n}$. Next, we define the privacy constraint considered in this work.
Definition 1 ($(\epsilon,\delta)$-DP in DME)
Let $\mathcal{G}_{i}=\{\{M(\mathbf{x}_{j})\}_{j\in[1:n]\setminus\{i\}},\{\mathbf{x}_{j},\mathbf{Z}_{j},\mathcal{I}_{j}\}_{j\in\mathcal{C}}\}$ denote the information observed by the server from all users except user $i$, for any non-colluding user $i\in[1:n]\setminus\mathcal{C}$. For a given $\epsilon\geq 0$ and $\delta\in(0,1)$, a DP-DME scheme ensures $(\epsilon,\delta)$-DP if the following is satisfied.
$\mathbb{P}(M(\mathbf{x}_{i})\in\mathcal{A}|\mathcal{G}_{i})\leq e^{\epsilon}\mathbb{P}(M(\mathbf{x}_{i}^{\prime})\in\mathcal{A}|\mathcal{G}_{i})+\delta,\quad\forall i\in[1:n]\setminus\mathcal{C}$ (3)
for any fixed $\mathbf{x}_{j}\in\mathbb{B}^{d}$, $j\neq i$, $\forall\mathbf{x}_{i},\mathbf{x}_{i}^{\prime}\in\mathbb{B}^{d}$ and $\forall\mathcal{A}\subset\mathbb{R}^{d}$ in the Borel $\sigma$-field.
As the privacy constraint in (3) must be satisfied for any fixed $\mathbf{x}_{j}$,$j\neq i$ for each $i$, Definition 1 implies that $(\epsilon,\delta)$-DP of $\mathbf{x}_{i}$ is satisfied even if the server has access to all $\mathbf{x}_{j}$, $j\neq i$, in addition to $M(\mathbf{x}_{j})$,$\forall j$, and all the information from the colluding users. The privacy constraint in Definition 1 encapsulates the fundamental principles of the original differential privacy (DP) framework as follows. Let $\mathcal{D}=\{\mathbf{x}\}_{j\neq i}\cup\mathbf{x}_{i}$ and $\mathcal{D}^{\prime}=\{\mathbf{x}\}_{j\neq i}\cup\mathbf{x}_{i}^{\prime}$ be two neighboring datasets that only differ in the vector of user $i$ for any $i\in[1:n]\setminus\mathcal{C}$. For a given dataset $\mathcal{D}$, the information that the server observes is denoted by $\mathcal{V}_{\mathcal{D}}=\{\{M(\mathbf{x}_{j})\}_{j=1}^{n},\{\mathbf{x}_{j},\mathbf{Z}_{j},\mathcal{I}_{j}\}_{j\in\mathcal{C}}\}\in\mathcal{J}$, where $\mathcal{J}$ is the domain of $\mathcal{V}_{\mathcal{D}}$. The original DP framework requires that, for any given $(\epsilon,\delta)$,
$\sup_{\mathcal{D}\sim\mathcal{D^{\prime}}}\mathbb{P}(\mathcal{V}_{\mathcal{D}}\in\mathcal{Y})\leq e^{\epsilon}\mathbb{P}(\mathcal{V}_{\mathcal{D}^{\prime}}\in\mathcal{Y})+\delta$ (4)
must be satisfied for all $\mathcal{Y}\subset\mathcal{J}$. By interpreting (4) as $\mathbb{P}\left(\ln|\frac{\mathbb{P}(\mathcal{V}_{\mathcal{D}}\in\mathcal{Y})}{\mathbb{P}(\mathcal{V}_{\mathcal{D^{\prime}}}\in\mathcal{Y})}|\right)\leq\epsilon)\geq 1-\delta$, for all $\mathcal{D}$,$\mathcal{D}^{\prime}$ [38], we see that (4) is equivalent to (3), when satisfied $\forall\mathbf{x}_{i},\mathbf{x}_{i}^{\prime}$, for all fixed $\{\mathbf{x}_{j}\}_{j\neq i}$ and for all $i\in[1:n]\setminus\mathcal{C}$. The privacy constraint in Definition 1 is a generalization of the privacy constraints used in LDP and distributed DP with SecAgg, which can be used to investigate the privacy guarantees of an adversarial DME setting. The proofs LDP and distributed DP with SecAgg being special cases of the privacy constraint in Definition 1 are given in appendix A.
Definition 2
An $(n,t,c)$-private DME scheme characterized by the encoding and decoding functions in (1)-(2) is defined by the joint distribution of $(\mathbf{Z}_{1},\dotsc,\mathbf{Z}_{n})$ that satisfy Definition 1 and the linear decoding functions $d_{\mathcal{U}}:\mathbb{R}^{d|\mathcal{U}|}\rightarrow\mathbb{R}^{d}$, for each subset of responding but non-colluding users $\mathcal{U}\subseteq[1:n]$.
For a given $(\epsilon,\delta)$-DP constraint, the accuracy of an $(n,t,c)$-private DME scheme is measured by the MSE between the mean estimate at the server $\hat{\mathbf{S}}_{\mathcal{U}}$ and the true mean $\mathbf{S}_{\mathcal{U}}=\frac{1}{|\mathcal{U}|}\sum_{i\in{\mathcal{U}}}\mathbf{x}_{i}$.
Definition 3 (MSE of $(n,t,c)$-private DME)
Consider a DME scheme satisfying $(\epsilon,\delta)$-DP with a given joint distribution $\mathcal{D}_{Z}$ of $(\mathbf{Z}_{1},\dotsc,\mathbf{Z}_{n})$ and a decoder $d_{\mathcal{U}}$. For any $\mathcal{F}$ and $\mathcal{C}$, the MSE is defined as,
$\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\mathcal{D}_{Z},\bm{\alpha})\triangleq\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\mathbb{E}[\|\hat{\mathbf{S}}_{\mathcal{U}}-\mathbf{S}_{\mathcal{U}}\|_{2}^{2}]$ (5)
where $\bm{\alpha}=[\alpha_{j_{1}},\dotsc,\alpha_{j_{|\mathcal{U}|}}]^{T}$. The expectation is over $\mathcal{D}_{Z}$.
The goal of this work is to characterize the minimum MSE of a given $(n,t,c)$-private DME setting with a given $(\epsilon,\delta)$-DP constraint. We navigate the trade-off between MSE and DP constraints (3) by optimizing the distribution of the privacy mechanism $\mathcal{D}_{Z}$ and the linear decoder at the server:***The order of optimizations in (6) is explained as follows. The privacy mechanism $\mathcal{D}_{Z}$ is determined prior to the mean estimation step in (2), and is fixed irrespective of the number of responding users at a given time. Therefore, the problem is formulated to optimize $\mathcal{D}_{Z}$ for the worst case dropouts and colluding users. The decoder on the other hand is used at the time of estimation as shown in (2), and is optimized based on the number of responding users for any fixed $\mathcal{D}_{Z}$. This allows the decoder to make use of the information on the set of responding users at any given time.
$\mathrm{MMSE}\triangleq\inf_{\mathcal{D}_{Z}}\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n\end{subarray}}\inf_{\bm{\alpha}}\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\mathbb{E}[\|\hat{\mathbf{S}}_{\mathcal{U}}-\mathbf{S}_{\mathcal{U}}\|_{2}^{2}]$ (6)
3 Correlated Gaussian Mechanism for DP-DME: A Geometric Interpretation
In this section, we provide a geometric interpretation of the DP-DME problem, where we show how carefully tuned (anti) correlated noise can significantly outperform DME with independent noise (LDP) while ensuring the same level of privacy, in the absence of dropouts and collusions. This also presents a re-interpretation of one of the core ideas of the SecAgg scheme - but over $\mathbb{R}^{d}$ in the DP framework rather than over finite fields. Building up on this idea, we analyze the privacy-utility trade-offs with the presence of dropouts and/or collusions, and show that carefully tuned correlated noise continues to provide superior performance to independent noise (see Section 5 for the results).
Consider a simple two-user setting with no dropouts and no colluding users, where the private vectors of the two users are given by $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, with $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathbb{S}^{d-1}$. User $i$ sends $M(\mathbf{x}_{i})=\mathbf{x}_{i}+\mathbf{Z}_{i}$ to the server, where $\mathbf{Z}_{i}\sim\mathcal{N}(\mathbf{0}_{d},\sigma^{2}\mathsf{I}_{d})$ for $i=1,2$. The elements of $\mathbf{Z}_{1}$ and $\mathbf{Z}_{2}$ are correlated as,
$\mathbb{E}[Z_{1,j}Z_{2,k}]=\begin{cases}\rho\sigma^{2},&j=k,\quad j,k\in\{1,\dotsc,d\}\\ 0,&j\neq k,\quad j,k\in\{1,\dotsc,d\}.\end{cases}$ (7)
[…]
Next, we geometrically illustrate the privacy constraint in (3). For this example, (3) simplifies to $\mathbb{P}(\mathbf{x}_{i}+\mathbf{Z}_{i}\in\mathcal{A}|\mathbf{Z}_{j}=\mathbf{y}_{j}-\mathbf{x}_{j})\leq e^{\epsilon}\mathbb{P}(\mathbf{x}_{i}^{\prime}+\mathbf{Z}_{i}\in\mathcal{A}|\mathbf{Z}_{j}=\mathbf{y}_{j}-\mathbf{x}_{j})+\delta$, for each $i\neq j$, for any fixed $\mathbf{x}_{j},\mathbf{y}_{j}$ and $\forall\mathbf{x}_{i},\mathbf{x}_{i}^{\prime}$. Let $\bm{Z}_{i}^{\parallel}:=\langle\bm{Z}_{i},\bm{Z}_{j}\rangle\bm{Z}_{j}/|\bm{Z}_{j}|^{2}=(\cos\theta)\bm{Z}_{j}$ and $\bm{Z}_{i}^{\bot}:=\bm{Z}_{i}-\bm{Z}_{j}^{\parallel}$ be the components of $\bm{Z}_{i}$ in $\mathcal{H}$, that are parallel and orthogonal to $\bm{Z}_{j}$, respectively, for $j\neq i$. The privacy constraint on $\bm{x}_{i}$ essentially imposes the lower bound: $|\bm{Z}_{i}^{\bot}|\geq\hat{\sigma}_{\epsilon,\delta}$, where $\hat{\sigma}_{\epsilon,\delta}$ is a constant that depends on the given privacy parameters and ||$\cdot$|| ⋅ | denotes the magnitude of a vector in $\mathcal{H}$ (See Appendix F for a rigorous proof). This is illustrated in Fig. 2, along with the vectors $\bm{Z}_{1}+\bm{Z}_{2}$ in red, that represent the MSEs corresponding to 1) $\rho<0$, 2) $\rho=0$, and 3) $\rho>0$. As we assume Gaussian noise, $\rho=0$ corresponds to independent noise, and hence translates to classical LDP. If $\rho=0,$ then the minimum MSE is $|\bm{Z}_{1}+\bm{Z}_{2}|^{2}=E[||\bm{Z}_{1}+\bm{Z}_{2}||^{2}]=2\hat{\sigma}_{\epsilon,\delta}^{2}$ (this can also be observed geometrically from Fig. 2). However, notice that when $\rho<0$, certain components of $\bm{Z}_{1}$ and $\bm{Z}_{2}$ cancel out and results in a reduced $|\bm{Z}_{1}+\bm{Z}_{2}|$, compared to the other two cases. In the context of Fig. 2, it is required to optimize $|\bm{Z}_{i}|$ and $\theta$, i.e., $\sigma^{2}$ and $\rho$, to minimize $|\bm{Z}_{1}+\bm{Z}_{2}|$ such that $|\bm{Z}_{i}^{\bot}|=|\bm{Z}_{i}|\sin{\theta}\geq\hat{\sigma}_{\epsilon,\delta}$ for $i=1,2$. The optimum $\sigma^{2}$ and $\rho$ are given by $\sigma^{2}_{*}\to\infty$ and $\rho_{*}\to-1$. This result is explained as follows. As illustrated in Fig. 3, $|\mathbf{Z}_{1}+\mathbf{Z}_{2}|$ decreases as $|\mathbf{Z}_{i}|$ and $\theta$ increase while satisfying $|\mathbf{Z}_{i}|\sin\theta=\hat{\sigma}_{\epsilon,\delta}$ for privacy. In the limit, when $|\mathbf{Z}_{i}|\to\infty$ and $\theta\to\pi$,***Note that increasing $\theta$ beyond $\pi$ is not optimal as $\theta=\pi+\eta$ and $\theta=\pi-\eta$ correspond to the same setting. $\mathbf{Z}_{1}+\mathbf{Z}_{2}$ aligns perpendicular to $\mathbf{Z}_{2}$, and $|\mathbf{Z}_{1}+\mathbf{Z}_{2}|\to\hat{\sigma}_{\epsilon,\delta}$, which is the minimum achievable $|\mathbf{Z}_{1}+\mathbf{Z}_{2}|$ while ensuring $|\mathbf{Z}_{i}|\sin\theta=\hat{\sigma}_{\epsilon,\delta}$.The resulting minimum MSE is given by $\mathbb{E}[\|\mathbf{Z}_{1}+\mathbf{Z}_{2}\|^{2}]=|\mathbf{Z}_{1}+\mathbf{Z}_{2}|^{2}=\hat{\sigma}^{2}_{\epsilon,\delta}$.
Figure 2: The privacy constraint on $\mathbf{x}_{1}$, described by a lower bound on $|\mathbf{Z}_{1}^{\bot}|$, and the MSE, represented by $|\mathbf{Z}_{1}+\mathbf{Z}_{2}|$.
Figure 3: Variation of $|\mathbf{Z}_{1}+\mathbf{Z}_{2}|$ with changing $|\mathbf{Z}_{i}|$.
Next, we analyze the value of $\hat{\sigma}_{\epsilon,\delta}$ for any given $\epsilon$ and $\delta$. Let $\sigma_{\epsilon,\delta}=\inf_{\hat{\sigma}>0}\{\hat{\sigma};\Phi\left(\frac{1}{\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{2}\right)-e^{\epsilon}\Phi\left(-\frac{1}{\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{2}\right)\leq\delta\}$ for any given $\epsilon$ and $\delta$. The standard Gaussian mechanism in DP states that $\mathbf{Z}\sim\mathcal{N}(\mathbf{0}_{d},\sigma^{2}\mathsf{I}_{d})$ with $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$ ensures $\mathbb{P}(\mathbf{v}+\mathbf{Z}\in\mathcal{A})\leq e^{\epsilon}\mathbb{P}(\mathbf{v}^{\prime}+\mathbf{Z}\in\mathcal{A})+\delta$ for any $\mathbf{v},\mathbf{v}^{\prime}\in\mathbb{R}^{d}$ with $\|\mathbf{v}-\mathbf{v}^{\prime}\|\leq 2$, and $\forall\mathcal{A}\subset\mathbb{R}^{d}$. Note that this is the setting of CDP where $\mathbf{v}$ and $\mathbf{v}^{\prime}$ represent the summations of the users’ vectors with all vectors fixed except for one, which results in $\|\mathbf{v}-\mathbf{v}^{\prime}\|\leq 2$ as each vector is placed within the unit ball. Equivalently, for the case of CDP, the estimate of $\mathbf{v}$ given by $\mathbf{v}+\mathbf{Z}$ incurs an MSE that is lower bounded by,
$\mathbb{E}[\|\mathbf{Z}\|^{2}]\geq\sigma^{2}_{\epsilon,\delta}d$ (8)
to satisfy $(\epsilon,\delta)$-DP. Recall that when considering the privacy of $\mathbf{x}_{i}$ in the two user case, the effective noise added to $\mathbf{x}_{i}$ is quantified by $\mathbf{Z}_{i}^{\bot}$ as the rest of $\mathbf{Z}_{i}$ can be inferred by $\mathbf{Z}_{j}$, for $i,j\in\{1,2\}$, $j\neq i$. The privacy constraint in (3) reduces to $\mathbb{P}(\mathbf{x}_{i}+\mathbf{Z}_{i}^{\bot}\in\mathcal{A}^{\prime})\leq e^{\epsilon}\mathbb{P}(\mathbf{x}^{\prime}_{i}+\mathbf{Z}_{i}^{\bot}\in\mathcal{A}^{\prime})+\delta,\quad i=\{1,2\}$ for $i=\{1,2\}$, $\forall\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{B}^{d}$, and $\forall\mathcal{A}^{\prime}\subset\mathbb{R}^{d}$ in the Borel $\sigma$-algebra. This imposes a lower bound on the variance of $\mathbf{Z}_{i}^{\bot}$, and hence on the MSE resulted by the effective noise: $\mathbb{E}[\|\mathbf{Z}_{i}^{\bot}\|^{2}]=|\mathbf{Z}_{i}^{\bot}|^{2}\geq\sigma^{2}_{\epsilon,\delta}d$ to ensure $(\epsilon,\delta)$-DP based on the standard Gaussian mechanism, similar to (8). This implies $\hat{\sigma}_{\epsilon,\delta}=\sigma_{\epsilon,\delta}\sqrt{d}$.
Recall from (8) that the minimum MSE of DME with CDP is given by $\sigma^{2}_{\epsilon,\delta}d=\hat{\sigma}^{2}_{\epsilon,\delta}$. This (two-user) example shows that the same MSE can be achieved without the requirement of a trusted server by carefully choosing the parameters of the correlated privacy mechanism; notably, this is a factor of $2$ reduction as compared to LDP ($\rho=0$). The insights of Fig. 2 generalize for more than two users. Specifically, in Theorem 1 and Corollary 1, we show that the same MSE of $\sigma^{2}_{\epsilon,\delta}d$ is achieved for the general case of $n\geq 2$ users with carefully correlated privacy mechanisms, even with no trusted server. If the privacy mechanisms are not correlated, i.e., if $\mathbf{Z}_{i}\sim\mathcal{N}(\mathbf{0}_{d},\sigma^{2}\mathsf{I}_{d})$ are independent for $i\in[1:n]$, $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$ must hold for privacy and the estimation error is $\sum_{i=1}^{n}\mathbf{Z}_{i}$, which results in a minimum MSE of $\mathbb{E}[\|\sum_{i=1}^{n}\mathbf{Z}_{i}\|^{2}]=n\sigma^{2}_{\epsilon,\delta}d$, which is $n$ times larger than the MSE achieved by DME with CorDP-DME for the same level of privacy.
The correlated Gaussian scheme described above faces a significant shortcoming. If a user drops out, the residual noise of the remaining user/s cannot be canceled resulting in significantly high MSEs; this is because it uses noise with arbitrarily large variance (recall that the lowest possible MSE is achieved when $|\mathbf{Z}_{i}|^{2}=E[||\mathbf{Z}_{i}||^{2}]\to\infty$). In fact, SecAgg faces the same shortcoming. In SecAgg, users’ data is quantized to a finite field, and then a noise random variable that is uniformly distributed over the field elements is added to the quantized data and sent to the server. If a user drops out, this added noise cannot be canceled and the server input from the remaining users is statistically independent of the users’ data - which in effect results in a large MSE. SecAgg circumvents dropouts through additional rounds where remaining users’ share these non-canceled noise variables with the server enabling the server to cancel it***In fact, SecAgg is complex and requires even an additional third round in order it has to guard against data leaks and malicious behaviour in this second round as well.. However, our interpretation here opens the door to handling user dropouts much more directly, leading CorDP-DME protocol. In CorDP-DME we optimize the noise parameters—$\rho$ and $\sigma^{2}$—directly accounting for dropouts (up to a threshold). As our analysis demonstrates, this approach significantly improves upon LDP in terms of MSE, while avoiding additional rounds of communication. In fact, in Section 5, we show that for any dropout threshold, CorDP-DME outperforms LDP with independent Gaussian noise.
4 Proposed Approach
In this paper, we study the case where $\mathcal{D}_{Z}$ is multivariate Gaussian. Specifically, we choose $\mathbf{Z}_{i}\sim\mathcal{N}(\mathbf{0}_{d},\sigma^{2}\mathsf{I}_{d})$, $i\in[1:n]$, with the all zeros vector of size $d\times 1$ denoted by $\mathbf{0}_{d}$ and the identity matrix of size $d\times d$ denoted by $\mathsf{I}_{d}$. The $k$th element of $\mathbf{Z}_{i}$ is denoted by $Z_{i,k}$ for $i\in[1:n]$ and $k\in[1:d]$. $Z_{i,k}$’s are distributed as, $[Z_{1,k},\dotsc,Z_{n,k}]^{T}\sim\mathcal{N}(\mathbf{0}_{n},\Sigma)$, for $k\in[1:d]$, where $\Sigma_{i,i}=\sigma^{2}$ and $\Sigma_{i,j}=r$ for $i,j\in[1:n]$, $i\neq j$, i.e.,
$\begin{pmatrix}Z_{1,k}\\ Z_{2,k}\\ \vdots\\ Z_{n,k}\end{pmatrix}\sim\mathcal{N}\left(\mathbf{0}_{n},\begin{pmatrix}\sigma^{2}&r&\dotsc&r\\ r&\sigma^{2}&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots\\ r&r&\dotsc&\sigma^{2}\end{pmatrix}\right),\quad k\in[1:d]$ (9)
The correlation coefficient is denoted by $\rho=\frac{r}{\sigma^{2}}$. Moreover, we let $\mathbb{E}[Z_{i,k}Z_{j,k^{\prime}}]=0$, $\forall i,j$, $\forall k\neq k^{\prime}$. We denote this class of distributions as $\mathcal{D}_{Z}^{G}$. Next, we present the noise generation protocol used to achieve the exact covariance structure among the privacy mechanisms of the distributed users, specified by $\Sigma$. Based on the values of $n,t,c,\epsilon,\delta$ considered in the system, the optimum noise parameters $\sigma^{2}_{*}$ and $r_{*}$ are calculated first (see Section 5 for optimum values). Then, the following steps generate the explicit noise terms with the required correlation structure.
  1. 1.
    Each pair of users $(i,j)$, $i,j\in[1:n]$, $i\neq j$ generates a pairwise random seed using the Diffie-Hellman key exchange.
  2. 2.
    Using the common seed, each pair of users $(i,j)$ samples the same random vector $\mathbf{S}_{i,j}\in\mathbb{R}^{d}$ from $\mathcal{N}(\mathbf{0}_{d},-r_{*}\mathsf{I}_{d})$,***We use use negative correlation among the privacy mechanisms of different users. Hence, $-r_{*}>0$. where $\mathbf{S}_{i,j}=\mathbf{S}_{j,i}$, and $\mathbf{S}_{i,j}$, $\mathbf{S}_{i^{\prime},j^{\prime}}$ are independent for any $(i,j)\neq(i^{\prime},j^{\prime})$.
  3. 3.
    Each user $i$ independently generates another noise variable $\mathbf{N}_{i}\sim\mathcal{N}(\mathbf{0}_{d},\sigma^{2}_{*}+r_{*}(n-1)\mathsf{I}_{d})$.
  4. 4.
    The combined noise added to $\mathbf{x}_{i}$ $i\in[1:n]$ is: $\mathbf{Z}_{i}=\sum_{j=1,j<i}^{n}\mathbf{S}_{i,j}-\sum_{j=1,j>i}^{n}\mathbf{S}_{i,j}+\mathbf{N}_{i}$. Then user $i$, $i\in[1:n]$ sends $M(\mathbf{x}_{i})=\mathbf{x}_{i}+\mathbf{Z}_{i}$ to the server.
5 Main Results
In this section, we provide the solution to (6) while satisfying (3), for the class of multivariate Gaussian distributions $\mathcal{D}_{Z}^{G}$ specified in Section 4. The following notation is used throughout the paper.
$\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},\rho,\bm{\alpha})=\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\mathbb{E}\left[\left\|\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\alpha_{i}(\mathbf{x}_{i}+\mathbf{Z}_{i})-\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\mathbf{x}_{i}\right\|^{2}_{2}\right]$ (10)
Whenever $t=n$ (no dropouts) or $c=0$ (no collusion), we drop the $\mathcal{F}$ and $\mathcal{C}$ terms in the MSE notation in (10). In Proposition 1, we state the optimum decoder, $\bm{\alpha}^{*}$, (with $\alpha_{i}^{*}$ as its $i$th component) that minimizes the MSE for any given $\mathcal{F},\mathcal{C},\sigma^{2}$ and $\rho$.
Proposition 1 (Optimum decoder)
For any fixed $\mathcal{F},\mathcal{C}\subset[1:n]$ satisfying $|\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n$, and for any $\sigma^{2}$, $\rho$, the optimum decoder is given by,
$\bm{\alpha}^{*}=\arg\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})=\frac{1}{1+\frac{d\sigma^{2}}{|\mathcal{U}|}(1+\rho(|\mathcal{F}|-1))}\mathbf{1}_{|\mathcal{U}|}$ (11)
Theorem 1 characterizes the MMSE in (6) for $\mathcal{D}_{Z}=\mathcal{D}_{Z}^{G}$ while ensuring the privacy constraint in Definition 1, for any $t$ and $c$. Note that finding the optimum $\mathcal{D}_{Z}^{G}$ in (6) is equivalent to optimizing $r$ (or equivalently $\rho$) and $\sigma^{2}$ that minimizes the MSE while satisfying the privacy constraint in Definition 1, based on the structure of $\mathcal{D}_{Z}^{G}$ specified in Section 4.
Theorem 1
For any given $\epsilon>0$, $\delta\in(0,1)$, $t$ and $c$, the optimum $\sigma^{2}$ and $\rho$ that solves (6) for $\mathcal{D}_{Z}=\mathcal{D}_{Z}^{G}$ while satisfying (3) is characterized by,
$(\sigma^{2}_{*},\rho_{*})=\arg\min_{\sigma^{2},\rho}\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n\end{subarray}}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},\rho,\bm{\alpha})$
$\sigma^{2}_{*}$ =$=\begin{cases}\infty,&t=n\\ \frac{\sigma^{2}_{\epsilon,\delta}(n^{2}-2n-cn+2)}{(n-c)^{2}}+\frac{\sigma^{2}_{\epsilon,\delta}(n-c-1)(n+c-2nc+t(n+c-2))}{(n-c)^{2}\sqrt{(t-c)(n-t)(n-c-1)}},&c<t<n.\end{cases}$ (12)
$\rho_{*}$ =$=\begin{cases}-\frac{\sigma^{2}_{*}-\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2)},&c=1\\ \frac{-(n-2)\left(1-\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}}\right)-c}{2(n-1)(c-1)}+\frac{\sqrt{\left((n-2)\left(1-\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}}\right)-c\right)^{2}+4(n-c-1)\left(1-\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}}\right)}}{2(n-1)(c-1)},&c\neq 1\end{cases}$ (13)
with $\sigma_{\epsilon,\delta}\!=\!\inf_{\hat{\sigma}>0}\{\hat{\sigma};\Phi\!\left(\frac{1}{\hat{\sigma}}\!-\!\frac{\epsilon\hat{\sigma}}{2}\right)-e^{\epsilon}\Phi\!\left(-\frac{1}{\hat{\sigma}}\!-\!\frac{\epsilon\hat{\sigma}}{2}\right)\leq\delta\}$, where $\Phi$ is the standard Gaussian CDF. The resulting minimum MSE is given by,
$\min_{\sigma^{2},\rho}\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\geq|\mathcal{F}|\leq n\end{subarray}}$ $\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},\rho,\bm{\alpha})=\left(1+\frac{(t-c)/d}{\sigma^{2}_{*}+\rho_{*}(t-1)}\right)^{-1}$ (14)
For a fixed $\epsilon$ and $\delta$, the value of $\sigma^{2}_{\epsilon,\delta}$ in Theorem 1 is the minimum variance of the Gaussian noise added to achieve $(\epsilon,\delta)$-DP in the standard Gaussian mechanism with an $L_{2}$ sensitivity of $2$. For any $\epsilon,\delta,t$ and $c$, Theorem 1 shows that $\rho_{*}\leq 0$ always holds. This implies that the (anti) correlated Gaussian mechanism outperforms (or performs equally when $c=t-1$) the independent Gaussian mechanism for any $\epsilon,\delta,t,c$. Next, we upper bound $\sigma^{2}_{\epsilon,\delta}$ using the results from [39, 40, 8] to better interpret the dependency of the MSE on $\epsilon$ and $\delta$. Generally, we assume that $\delta=10^{-5}$.
Proposition 2
The following upper bounds hold for $\sigma^{2}_{\epsilon,\delta}$, where $\eta=1+2\sqrt{\ln\frac{1}{2\delta}}$ for $\delta\in(0,0.05]$ and $\eta=1+2\sqrt{\ln 10}$ for $\delta\in(0.05,1)$.
$\sigma^{2}_{\epsilon,\delta}$$\leq\begin{cases}\frac{8\ln(1.25/\delta)}{\epsilon^{2}},&\epsilon,\delta\in(0,1)\\ \frac{2\eta^{2}}{\epsilon},&\epsilon\geq 1,\delta\in(0,1)\end{cases}$ (15)
In Corollaries 13, we consider special cases of Theorem 1, and provide simplified MSE results using the bounds in Proposition 2.
Corollary 1 (Without collusion, Without dropouts)
For $t=n$, $c=0$, the minimum MSE is given by,
$\min_{\sigma^{2},\rho,\bm{\alpha}}\mathrm{MSE}(n,\sigma^{2},\rho,\bm{\alpha})$ =$=\begin{cases}O\left(\frac{d}{n^{2}}\frac{\ln(1/\delta)}{\min\{\epsilon,\epsilon^{2}\}}\right),&\text{if }n^{2}>>d\\ O(1),&otherwise.\end{cases}$
Corollary 2 (Without collusion, with dropouts)
For $t<n$ and $c=0$, the minimum MSE is given by,
$\min_{\sigma^{2},\rho}\max_{\begin{subarray}{c}\mathcal{F}\subset[1:n]\\ t\leq|\mathcal{F}|<n\end{subarray}}$ $\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\sigma^{2},\rho,\bm{\alpha})=\begin{cases}O\left(\frac{d(n-t)}{tn}\frac{\ln(1/\delta)}{\min\{\epsilon,\epsilon^{2}\}}\right),&\text{if }\frac{n^{2}}{n-t}>>d\\ O(1),&otherwise.\end{cases}$ (16)
Corollary 3 (With collusion, without dropouts)
For $t=n$ and $c>0$ the minimum MSE is given by,
$\min_{\sigma^{2},\rho}\quad\max_{\begin{subarray}{c}\mathcal{C}\subset[1:n]\\ 0\leq|\mathcal{C}|\leq c\end{subarray}}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{C},\sigma^{2},\rho,\bm{\alpha})=\begin{cases}O\left(\frac{d}{(n-c)^{2}}\frac{\ln(1/\delta)}{\min\{\epsilon,\epsilon^{2}\}}\right),&\text{if }(n-c)^{2}>>d\\ O(1),&otherwise.\end{cases}$
For the case of no dropouts (with or without colluding users) CorDP-DME achieves the same order of MSE as CDP (and DP-DME with SecAgg) for a fixed $\delta$, as shown in Corollaries 1 and 3. Note that for $|\mathcal{U}|$ non-colluding and responding users, CorDP-DME achieves a reduction in the MSE by a factor of $O(|\mathcal{U}|)$, compared to LDP. For the case of dropouts, CorDP-DME achieves MSEs that lie in between LDP and CDP (Corollary 2). If the maximum number of dropouts is $O(n^{p})$ for any $p<1$, Corollary 2 implies that the minimum MSE is $O(d\sigma^{2}_{\epsilon,\delta}/(tn^{1-p}))$, which has a scaling advantage over the case with independent noise, i.e., $\rho=0$, with a minimum MSE of $O(d\sigma^{2}_{\epsilon,\delta}/t)$.
The optimum decoder in Proposition 1 results in a biased estimate of the mean. By fixing the decoder to be $\bm{\alpha}=\mathbf{1}_{|\mathcal{U}|}$, we can optimize the noise distribution and obtain an unbiased estimate with MSEs that lie in between the corresponding MSEs of DME with LDP and CDP. Proposition 3 characterizes the optimum noise distribution and the resulting MSE (unbiased) for the case where $\bm{\alpha}=\mathbf{1}_{|\mathcal{U}|}$.
Proposition 3
For any $\epsilon>0$ and $\delta\in(0,1)$, the $r_{*}$ and $\sigma^{2}_{*}$ given in (13) and (12) satisfies,
$(\sigma^{2}_{*},\rho_{*})=\arg\min_{\sigma^{2},\rho}\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n\end{subarray}}\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\mathbb{E}\left[\left\|\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}(\mathbf{x}_{i}+\mathbf{Z}_{i})-\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\mathbf{x}_{i}\right\|^{2}_{2}\right],$ (17)
and the resulting MSE is given by $\frac{d(\sigma^{2}_{*}+r_{*}(t-1))}{t-c}$.
The simplified MSEs of the unbiased case corresponding to the settings in Corollaries 1-3 are given by $O\left(\frac{d}{n^{2}}\frac{\ln(1/\delta)}{\min\{\epsilon,\epsilon^{2}\}}\right)$, $O\left(\frac{d(n-t)}{tn}\frac{\ln(1/\delta)}{\min\{\epsilon,\epsilon^{2}\}}\right)$ and $O\left(\frac{d}{(n-c)^{2}}\frac{\ln(1/\delta)}{\min\{\epsilon,\epsilon^{2}\}}\right)$, respectively.
6 Experiments
We implement CorDP-DME for specific values of $\epsilon,\delta,n,t,c$, and compare the privacy-utility trade-offs against DME with LDP and CDP (with the Gaussian mechanism).***The experiments are performed on commodity hardware/laptop. As the two baselines correspond to unbiased estimates, we compare both biased (see Theorem 1) and unbiased (see Proposition 3) versions of CorDP-DME with LDP and CDP. Fig. 4 shows the privacy-utility trade-offs of different DP-DME cases corresponding to $n=100$, $\delta=10^{-5}$ and $d=20$. Recall that $t$ and $c$ are the thresholds on the minimum responding users and maximum colluding users, respectively. CorDP-DME coincides with CDP when no dropouts or colluding users are considered (left). When dropouts and/or colluding users are considered, the MSE of CorDP-DME lies in between the MSEs of (Gaussian) LDP and CDP (middle/ right). We also show that for many cases (i.e., $t=80$, $c=20$ on the right), CorDP-DME outperforms PrivUnitG [20], which is an approximation of PrivUnit [22], that is proven to be the optimum among all LDP mechanisms that result in unbiased estimates.
Figure 4: Privacy-utility trade-offs with no dropouts and no colluding users (left), either dropouts or colluding users (middle), both dropouts and colluding users (right).
7 Conclusions and Limitations
In this work, we present CorDP-DME, a novel differentially private DME protocol that uses correlated Gaussian noise to achieve a favorable balance between utility, resilience to dropouts, and robustness against colluding users. CorDP-DME spans the spectrum between DME with LDP, which provides strong resilience but poor utility, and SecAgg-based approaches, which achieve high utility but have limited resilience to dropouts and collusion. A key insight is that carefully tuned (anti) correlated noise can significantly improve utility compared to independent noise mechanisms, even in adversarial settings with dropouts and collusions. One limitation of CorDP-DME in its current form is that it operates in the continuous domain using Gaussian noise. While this allows for rigorous theoretical analysis, it may present challenges for direct practical implementation such as communication inefficiencies and vulnerabilities to floating point attacks [41, 42]. An important direction for future work is to develop a discrete or quantized variants of CorDP-DME, as well as extending the protocol to non-additive mechanisms.
In terms of theoretical extensions, CorDP-DME opens up several intriguing questions. Is Gaussian noise optimal in this correlated noise setting, or can other noise distributions yield further improvements? How can the noise correlations be optimally designed for different network topologies? Studying these questions may yield further insights and advances.
Overall, CorDP-DME takes an important step towards bridging the gap between LDP and SecAgg-based protocols. The correlated Gaussian noise framework provides a foundation for the future design of DP mechanisms that can achieve high accuracy while maintaining robustness in distributed settings. Our information-theoretic analysis of CorDP-DME will inspire further correlated noise mechanisms for distributed privacy-preserving learning.
Appendix A Generalized Privacy Definition
In this section, we show that the privacy definitions used in DME with LDP and SecAgg based distributed DP are special cases of the proposed privacy constraint in Definition 1.
DME with LDP:
Recall that each $\mathcal{G}_{i}$ in Definition 1 contains information of all users except user $i$. Therefore, the random variables in $\mathcal{G}_{i}$ are independent of $M(\mathbf{x}_{i})$, for all $i\in[1:n]\setminus\mathcal{C}$ as DME with LDP utilizes independent privacy mechanisms among users. Thus, (3) directly simplifies to,
$\mathbb{P}(M(\mathbf{x}_{i})\in\mathcal{A})\leq e^{\epsilon}\mathbb{P}(M(\mathbf{x}_{i}^{\prime})\in\mathcal{A})+\delta,\quad\forall i\in[1:n]\setminus\mathcal{C},$ (18)
$\forall\mathbf{x}_{i},\mathbf{x}_{i}^{\prime}\in\mathbb{B}^{d}$ and $\forall\mathcal{A}\subset\mathbb{R}^{d}$ in the Borel $\sigma$-field, which is the privacy constraint in DME with LDP.
SecAgg based DME with distributed DP:
To prove that the privacy constraint used in SecAgg based DME with distributed DP is a special case of the privacy constraint in Definition 1, we first consider what is transmitted and received by the users and the server, respectively.
user $i$rightarrow server: $\bar{M}(\mathbf{x}_{i})=M(\mathbf{x}_{i})+\sum_{j<i}\mathbf{\bar{S}}_{j,i}-\sum_{j>i}\mathbf{\bar{S}}_{i,j},\quad\forall i\in[1:n]$ (19)
where $\mathbf{\bar{S}}_{i,j}$ are uniformly distributed random variables from a finite field $\mathbb{F}_{p}$. Considering no dropouts and no colluding users, the privacy constraint in Definition 1 implies,
$\mathbb{P}(\bar{M}(\bm{x}_{i})\in\mathcal{A}|\bar{M}(\mathbf{x}_{j})=y_{j},\forall j\neq i)\leq e^{\epsilon}\mathbb{P}(\bar{M}(\bm{x}_{i}^{\prime})\in\mathcal{A}|\bar{M}(\mathbf{x}_{j})=y_{j},\forall j\neq i)\!+\!\delta,\quad\forall i$ (20)
for any fixed $\mathbf{x}_{j}\in\mathbb{B}^{d}$, $j\neq i$, $\forall\mathbf{x}_{i},\mathbf{x}_{i}^{\prime}\in\mathbb{B}^{d}$ and $\forall\mathcal{A}\subset\mathbb{R}^{d}$ in the Borel $\sigma$-field. Let $\mathcal{A}^{\prime}=\{a^{\prime}:a^{\prime}=a+\sum_{j\neq i}y_{j},\forall a\in\mathcal{A}\}$.
$\mathbb{P}(\bar{M}(\bm{x}_{i})\in\mathcal{A}|\bar{M}(\mathbf{x}_{j})=y_{j},\forall j\neq i)$
$=\frac{\mathbb{P}(\bar{M}(\bm{x}_{i})\in\mathcal{A},\sum_{j=1}^{n}\bar{M}(\mathbf{x}_{j})\in\mathcal{A}^{\prime},\{\bar{M}(\mathbf{(}x)_{j})=y_{j}\}_{j\neq i})}{\mathbb{P}(\{\bar{M}(\mathbf{(}x)_{j})=y_{j}\}_{j\neq i})}$ (21)
$=\mathbb{P}(\bar{M}(\bm{x}_{i})\in\mathcal{A}|\sum_{j=1}^{n}\bar{M}(\mathbf{x}_{j})\in\mathcal{A}^{\prime},\{\bar{M}(\mathbf{(}x)_{j})=y_{j}\}_{j\neq i})$
×$\quad\qquad\times\mathbb{P}(\sum_{j=1}^{n}\bar{M}(\mathbf{x}_{j})\in\mathcal{A}^{\prime}|\{\bar{M}(\mathbf{(}x)_{j})=y_{j}\}_{j\neq i}))$ (22)
$=1\times\frac{\mathbb{P}(\sum_{j=1}^{n}\bar{M}(\mathbf{x}_{j})\in\mathcal{A}^{\prime},\{\bar{M}(\mathbf{(}x)_{j})=y_{j}\}_{j\neq i})}{\mathbb{P}(\{\bar{M}(\mathbf{(}x)_{j})=y_{j}\}_{j\neq i})}$ (23)
$=\frac{\mathbb{P}(\sum_{j=1}^{n}M(\mathbf{x}_{j})\in\mathcal{A}^{\prime})\mathbb{P}(\{\bar{M}(\mathbf{(}x)_{j})=y_{j}\}_{j\neq i})}{\mathbb{P}(\{\bar{M}(\mathbf{(}x)_{j})=y_{j}\}_{j\neq i})}=\mathbb{P}(\sum_{j=1}^{n}M(\mathbf{x}_{j})\in\mathcal{A}^{\prime})$ (24)
where we use $\sum_{i=1}^{n}\bar{M}(\mathbf{x}_{i})=\sum_{i=1}^{n}M(\mathbf{x}_{i})$ and Shannon’s one-time-pad theorem to derive the last two steps. Substituting (24) in (20) (with all $\mathbf{x}_{j}$, $j\neq i$ fixed and $\mathbf{x}_{i}$ on the LHS and $\mathbf{x}_{i}^{\prime}$ on the RHS) gives the privacy constraint in distributed DP with SecAgg.
Appendix B Proof of Proposition 1
Proposition 1 restated:
For any fixed $\mathcal{F},\mathcal{C}\subset[1:n]$ satisfying $|\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n$, and for any $\sigma^{2}$, $r$, the optimum decoder is given by,
$\bm{\alpha}^{*}=\arg\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})=\frac{1}{1+\frac{d}{|\mathcal{U}|}(\sigma^{2}+r(|\mathcal{F}|-1))}\mathbf{1}_{|\mathcal{U}|}$ (25)
Proof:  Let $\mathcal{U}=\mathcal{F}\setminus\mathcal{C}$, and let $x_{i,k}$ and $Z_{i,k}$ denote the $k$th element of $\mathbf{x}_{i}$ and $\mathbf{Z}_{i}$, respectively. Then,
$\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})$ $=\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\mathbb{E}\left[\left\|\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\alpha_{i}(\mathbf{x}_{i}+\mathbf{Z}_{i})-\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\mathbf{x}_{i}\right\|^{2}\right]$ (26)
$=\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\sum_{k=1}^{d}\mathbb{E}\left[\left(\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\alpha_{i}({x}_{i,k}+Z_{i,k})-\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}{x}_{i,k}\right)^{2}\right]$ (27)
$=\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\frac{1}{|\mathcal{U}|^{2}}\sum_{k=1}^{d}\left(\bm{\alpha}^{T}A_{k}\bm{\alpha}-2\cdot\mathbf{1}^{T}\mathbf{x}^{[k]}\mathbf{x}^{[k]T}\bm{\alpha}+\mathbf{1}^{T}\mathbf{x}^{[k]}\mathbf{x}^{[k]T}\mathbf{1}\right)$ (28)
where $\mathbf{x}^{[k]}=[x_{j_{1},k},\dotsc,x_{j_{|\mathcal{U}|},k}]^{T}$ and $A_{k}=\mathbf{x}^{[k]}\mathbf{x}^{[k]T}+\Sigma$, where $\Sigma$ is the covariance matrix of $[Z_{1,k},\dotsc,Z_{|\mathcal{U}|,k}]^{T}$ for all $k\in[1:d]$. Define,
$f(\bm{\alpha},\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma)=\frac{1}{|\mathcal{U}|^{2}}\sum_{k=1}^{d}\left(\bm{\alpha}^{T}A_{k}\bm{\alpha}-2\cdot\mathbf{1}^{T}\mathbf{x}^{[k]}\mathbf{x}^{[k]T}\bm{\alpha}+\mathbf{1}^{T}\mathbf{x}^{[k]}\mathbf{x}^{[k]T}\mathbf{1}\right).$ (29)
For any fixed $\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}$ and $\Sigma$, $f(\bm{\alpha},\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma)$ is convex in $\bm{\alpha}$ as all $A_{k}$, $k\in[1:d]$ are positive definite. For any given $\bm{\alpha}$ in (26), as $\mathbf{Z}_{i}$, $i\in\mathcal{U}$ are i.i.d. and $\mathbf{x}_{i}$, $i\in\mathcal{U}$ are chosen from the same set $\mathbb{B}^{d}$, we have,
$\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})$ $=\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\Pi_{j}(\bm{\alpha})),\quad j\in[1:|\mathcal{U}|!]$ (30)
where $\Pi_{j}(\bm{\alpha})$ denotes the $j$th permutation of the elements of $\bm{\alpha}$. Therefore,
$\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})$ $=\frac{1}{|\mathcal{U}|!}\sum_{j=1}^{|\mathcal{U}|!}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\Pi_{j}(\bm{\alpha}))$ (31)
$=\frac{1}{|\mathcal{U}|!}\sum_{j=1}^{|\mathcal{U}|!}\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}f\left(\Pi_{j}(\bm{\alpha}),\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma\right)$ (32)
$\geq\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\frac{1}{|\mathcal{U}|!}\sum_{j=1}^{|\mathcal{U}|!}f\left(\Pi_{j}(\bm{\alpha}),\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma\right)$ (33)
$\geq\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}f\left(\frac{1}{|\tau|!}\sum_{j=1}^{|\tau|!}(\Pi_{j}(\bm{\alpha})),\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma\right)$ (34)
$=\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}f\left(\mathbf{\tilde{\alpha}},\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma\right)$ (35)
where (34) is due to the convexity of $f(\bm{\alpha},\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma)$ in $\bm{\alpha}$, and $\mathbf{\tilde{\alpha}}=\left(\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\alpha_{i}\right)\mathbf{1}_{|\mathcal{U}|}$. This implies that for any decoder $\bm{\alpha}$ with $\mathcal{U}$, $\sigma^{2}$ and $r$ fixed, there exists a $\mathbf{\tilde{\alpha}}=[\tilde{\alpha}_{j_{1}},\dotsc,\tilde{\alpha}_{j_{|\mathcal{U}|}}]^{T}$ such that $\tilde{\alpha}_{k}=\tilde{\alpha}_{\ell}$, $\forall k,\ell$ satisfying $\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})\geq\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\mathbf{\tilde{\alpha}})$. Therefore, the optimum decoder is of the form $\bm{\alpha}=\alpha\mathbf{1}_{|\mathcal{U}|}$, and (26) can be written as,
$\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\alpha)$
$=\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\mathbb{E}\left[\left\|\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\alpha(\mathbf{x}_{i}+\mathbf{Z}_{i})-\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\mathbf{x}_{i}\right\|^{2}\right]$ (36)
$=\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\frac{1}{|\mathcal{U}|^{2}}\sum_{k=1}^{d}\left((\alpha-1)^{2}\sum_{i\in\mathcal{U}}x_{i,k}^{2}+2(\alpha-1)^{2}\sum_{\begin{subarray}{c}i,j\in\mathcal{U}\\ i\neq j\end{subarray}}x_{i,k}x_{j,k}\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\left.+2\alpha^{2}\sum_{\begin{subarray}{c}i,j\in\mathcal{U}\\ i\neq j\end{subarray}}r+\alpha^{2}(\sigma^{2}+r|\mathcal{F}\cap\mathcal{C}|)|\mathcal{U}|\right)$ (37)
$=\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\frac{1}{|\mathcal{U}|^{2}}\left((\alpha-1)^{2}\sum_{i\in\mathcal{U}}\|\mathbf{x}_{i}\|^{2}+2(\alpha-1)^{2}\sum_{\begin{subarray}{c}i,j\in\mathcal{U}\\ i\neq j\end{subarray}}\mathbf{x}_{i}^{T}\mathbf{x}_{j}+d\alpha^{2}r|\mathcal{U}|(|\mathcal{U}|-1)\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\left.+d\alpha^{2}(\sigma^{2}+r|\mathcal{F}\cap\mathcal{C}|)|\mathcal{U}|\right)$ (38)
where (37) comes from $\text{Var}(\mathbf{Z}_{i})=-r(n-1-|\mathcal{F}\cap\mathcal{C}|)+\sigma^{2}+r(n-1)=\sigma^{2}+r|\mathcal{F}\cap\mathcal{C}|$. The worst case MSE is resulted when $\mathbf{x}_{i}^{T}\mathbf{x}_{j}=1$, for all $i,j\in\mathcal{U}$, that is, when $\mathbf{x}_{i}=\mathbf{x}_{j}$, $\forall i,j$, and $\mathbf{x}_{i}\in\mathbb{S}^{d-1}$, where $\mathbb{S}^{d-1}$ is the unit sphere in $\mathbb{R}^{d}$. Therefore,
$\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\alpha)$
$\quad=\frac{1}{|\mathcal{U}|^{2}}\left((\alpha-1)^{2}|\mathcal{U}|+(\alpha-1)^{2}|\mathcal{U}|(|\mathcal{U}|-1)+d\alpha^{2}r|\mathcal{U}|(|\mathcal{U}|-1)+d\alpha^{2}(\sigma^{2}+r|\mathcal{F}\cap\mathcal{C}|)|\mathcal{U}|\right)$ (39)
$\quad=(\alpha-1)^{2}+\frac{d\alpha^{2}}{|\mathcal{U}|}(\sigma^{2}+r(|\mathcal{F}|-1))$ (40)
as $|\mathcal{U}|+|\mathcal{F}\cap\mathcal{C}|=|\mathcal{F}|$. The optimum decoder for any fixed $\mathcal{U}$, $\sigma^{2}$ and $r$ is computed as,
$\frac{\partial\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\alpha)}{\partial\alpha}$ $=2(\alpha-1)+\frac{2d\alpha}{|\mathcal{U}|}(\sigma^{2}+r(|\mathcal{F}|-1))=0$ (41)
$\alpha^{*}$ $=\frac{1}{1+\frac{d}{|\mathcal{U}|}(\sigma^{2}+r(|\mathcal{F}|-1))}.$ (42)
The resulting MSE is obtained by substituting $\alpha^{*}$ in (40). $\blacksquare$
Appendix C Proof of Theorem 1
Theorem 1 restated:
For any given $\epsilon>0$, $\delta\in(0,1)$, $t$ and $c$, the optimum $\mathcal{D}_{Z}$ that solves (6) while satisfying (3) is characterized by,
$(\sigma^{2}_{*},\rho_{*})=\arg\min_{\sigma^{2},\rho}\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n\end{subarray}}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},\rho,\bm{\alpha})$
$\sigma^{2}_{*}$ =$=\begin{cases}\infty,&t=n\\ \frac{\sigma^{2}_{\epsilon,\delta}(n^{2}-2n-cn+2)}{(n-c)^{2}}+\frac{\sigma^{2}_{\epsilon,\delta}(n-c-1)(n+c-2nc+t(n+c-2))}{(n-c)^{2}\sqrt{(t-c)(n-t)(n-c-1)}},&c<t<n.\end{cases}$ (43)
$\rho_{*}$ =$=\begin{cases}-\frac{\sigma^{2}_{*}-\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2)},&c=1\\ \frac{-(n-2)\left(1-\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}}\right)-c}{2(n-1)(c-1)}+\frac{\sqrt{\left((n-2)\left(1-\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}}\right)-c\right)^{2}+4(n-c-1)\left(1-\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}}\right)}}{2(n-1)(c-1)},&c\neq 1\end{cases}$ (44)
with $\sigma_{\epsilon,\delta}\!=\!\inf_{\hat{\sigma}>0}\{\hat{\sigma};\Phi\!\left(\frac{1}{\hat{\sigma}}\!-\!\frac{\epsilon\hat{\sigma}}{2}\right)-e^{\epsilon}\Phi\!\left(-\frac{1}{\hat{\sigma}}\!-\!\frac{\epsilon\hat{\sigma}}{2}\right)\leq\delta\}$, where $\Phi$ is the standard Gaussian CDF. The resulting minimum MSE is given by,
$\min_{\sigma^{2},\rho}\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\geq|\mathcal{F}|\leq n\end{subarray}}$ $\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},\rho,\bm{\alpha})=\left(1+\frac{(t-c)/d}{\sigma^{2}_{*}+\rho_{*}(t-1)}\right)^{-1}$ (45)
Proof:  The proof consists of two main steps, namely, 1) optimizing the decoder at the server for any fixed joint distribution of $(\bm{Z}_{1},\dotsc,\bm{Z}_{n})$, i.e., any fixed $\sigma^{2}$ and $r$ (equivalently $\rho$), 2) determining the feasible regions of $\sigma^{2}$ and $r$ that satisfy the privacy constraint, 3) characterizing the overall minimum MSE by optimizing $\sigma^{2}$ and $r$. Step 1 is proved in Proposition 1. For step 2, note that the information available at the server increases with the number of responding users and the number of colluding users. Therefore, we consider the case where all $n$ users respond, and up to any $c$ users can collude with the server, to analyze the privacy constraint. Lemma  1 characterizes step 2. We analyze step 3 in two cases. In case 1, we assume no dropouts, i.e., $t=n$. In case 2, we assume that up to any $t<n$ users can dropout. The two cases are analyzed in Lemmas 2 and 3, respectively.
Let $D_{Z}^{G}(\sigma^{2},r)$ denote the multivariate Gaussian distribution (of $(\mathbf{Z}_{1},\dotsc,\mathbf{Z}_{n})$) with following properties. $\bm{Z}_{i}\sim\mathcal{N}(\bm{0}_{d},\sigma^{2}\mathsf{I}_{d})$, $i\in[1:n]$, with the all zeros vector of size $d\times 1$ denoted by $\bm{0}_{d}$ and the identity matrix of size $d\times d$ denoted by $\mathsf{I}_{d}$. The $k$th element of $\bm{Z}_{i}$ is denoted by $Z_{i,k}$ for $i\in[1:n]$ and $k\in[1:d]$. $Z_{i,k}$’s are allowed to be correlated as $\mathbb{E}[Z_{i,k}Z_{j,k}]=r$ for $\forall i\neq j$, $k\in[1:d]$. That is, $[Z_{1,k},\dotsc,Z_{n,k}]^{T}\sim\mathcal{N}(\bm{0}_{n},\Sigma)$ for $k\in[1:d]$ with $\Sigma_{i,i}=\sigma^{2}$ for $i\in[1:n]$ and $\Sigma_{i,j}=r$ for $i,j\in[1:n]$, $i\neq j$. Moreover, let $\mathbb{E}[Z_{i,k}Z_{j,k^{\prime}}]=0$, $\forall i,j$, $\forall k\neq k^{\prime}$.
Lemma 1 (Privacy Condition)
The Gaussian mechanism $\mathcal{D}_{Z}^{G}$ with given $\sigma^{2}$ and $r$ satisfies the privacy constraint in Definition 1 for any $\mathcal{F}$ and $\mathcal{C}$ so long as:
$r^{2}(n-1)(c-1)+r(\sigma^{2}(n+c-2)-\sigma^{2}_{\epsilon,\delta}(n-2))+\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})\geq 0,$ (46)
where $\sigma^{2}_{\epsilon,\delta}=\inf_{\hat{\sigma}>0}\{\hat{\sigma};\Phi\left(\frac{1}{\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{2}\right)-e^{\epsilon}\Phi\left(-\frac{1}{\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{2}\right)\leq\delta\}$. For any fixed $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$, the values of $r$ satisfying (46) are given by,
$r\geq\begin{cases}\frac{-\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2)},&c=1\\ \frac{-(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c}{2(n-1)(c-1)}+\frac{\sqrt{((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(n-1)(c-1)},&c>1\end{cases}$ (47)
and for $c=0$,
$r\in\left[a-b,a+b\right]$ (48)
where,
$a$ $=\frac{(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{2(n-1)}$ (49)
$b$ $=\frac{\sqrt{(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}+4(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(n-1)}.$ (50)
Proof:  [Proof of Lemma 1] Recall that the privacy mechanism of each user is given by $M(\mathbf{x}_{i})=\mathbf{x}_{i}+\mathbf{Z}_{i}$, $i\in[1:n]$. Note that for user $i$, the added noise can be written in the following form, based on our noise generation protocol in Section 4.
$\mathbf{Z}_{i}=\sum_{j<i}\mathbf{S}_{j,i}-\sum_{j>i}\mathbf{S}_{i,j}+N_{i},\quad i\in[1:n]$ (51)
where $\mathbf{S}_{i,j}\sim\mathcal{N}(\mathbf{0}_{d},-r\mathsf{I}_{d})$ and $\mathbf{N}_{i}\sim\mathcal{N}(\mathbf{0}_{d},(\sigma^{2}+r(n-1))\mathsf{I}_{d})$. To analyze the privacy constraint in Definition 1, the information available to the server from all users except for user $i$, $i\in[1:n]\setminus\mathcal{C}$ when any $\mathcal{C}\in[1:n]$, $|\mathcal{C}|\leq c$ users are allowed to collude with the server is given by,
$\mathcal{G}_{i}=\{\{M(\mathbf{x}_{j})\}_{j\in[1:n],j\neq i},\{\mathbf{x}_{k},\mathbf{N}_{k},\{\mathbf{S}_{k,j}\}_{\forall j}\}_{k\in\mathcal{C}}\}.$ (52)
Then, the privacy constraint in Definition 1 stated as
$\mathbb{P}(M(\mathbf{x}_{i})\in\mathcal{A}|\mathcal{G}_{i})\leq e^{\epsilon}\mathbb{P}(M(\mathbf{x}_{i}^{\prime})\in\mathcal{A}|\mathcal{G}_{i})+\delta,\quad\forall i\in[1:n]\setminus\mathcal{C},$ (53)
$\forall\mathbf{x}_{i},\mathbf{x}_{i}^{\prime}$, for any fixed $\mathbf{x}_{j}$, $j\neq i$ and $\forall\mathcal{A}$, simplifies to,
$\mathbb{P}(\hat{M}(\mathbf{x}_{i})\in\mathcal{A}|\hat{M}(\mathbf{x}_{j})=\mathbf{y}_{j},j\in[1:n],j\notin\mathcal{C},j\neq i)$
$\qquad\leq e^{\epsilon}\mathbb{P}(\hat{M}(\mathbf{x}_{i}^{\prime})\in\mathcal{A}|\hat{M}(\mathbf{x}_{j})=\mathbf{y}_{j},j\in[1:n],j\notin\mathcal{C},j\neq i)+\delta,\quad\forall i\in[1:n]\setminus\mathcal{C},$ (54)
where $\hat{M}(\mathbf{x}_{k})=\mathbf{x}_{k}+\sum_{j<k,j\notin\mathcal{C}}\mathbf{S}_{j,k}-\sum_{j>k,j\notin\mathcal{C}}\mathbf{S}_{k,j}+\mathbf{N}_{k}$. Let $\mathbf{Y}_{k}=\hat{M}(\mathbf{x}_{k})$, $\forall k\in[1:n]$. We first derive the conditional distribution of $\mathbf{Y}_{i}$ given $\mathbf{Y}_{j}$, $\forall j\notin\mathcal{C}\cup\{i\}$. Without loss of generality assume that $i=1$ and $\mathcal{C}=\{n-c+1,\dotsc,n\}$.
Claim 1
The conditional distribution of $\mathbf{Y}_{1}$, given $\mathbf{Y}_{j}$, $j\in[2:n-c]$ for any fixed $\mathbf{x}_{j}$, $j\in[2:n-c]$ is given by,
$\mathbf{Y}_{1}|\mathbf{Y}_{2}=\mathbf{y}_{2},\dotsc,\mathbf{Y}_{n-c}=\mathbf{y}_{n-c}\sim N(\tilde{\mathbf{\mu}},\tilde{\Sigma})$ (55)
where,
$\tilde{\mathbf{\mu}}$ $=\mathbf{x}_{1}+\left(r\mathbf{1}_{n-c-1}^{T}\begin{pmatrix}\sigma^{2}+rc&r&\dotsc&r\\ r&\sigma^{2}+rc&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots\\ r&r&\dotsc&\sigma^{2}+rc\end{pmatrix}_{n-c-1}^{-1}\otimes\mathsf{I}_{d}\right)\begin{pmatrix}\mathbf{y}_{2}-\mathbf{x}_{2}\\ \vdots\\ \mathbf{y}_{n-c}-\mathbf{x}_{n-c}\end{pmatrix}$ (56)
$\tilde{\Sigma}$ $=\left(\sigma^{2}+rc-r^{2}\mathbf{1}_{n-c-1}^{T}\begin{pmatrix}\sigma^{2}+rc&r&\dotsc&r\\ r&\sigma^{2}+rc&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots\\ r&r&\dotsc&\sigma^{2}+rc\end{pmatrix}_{(n-c-1)\times(n-c-1)}^{-1}\!\!\!\!\!\!\!\!\mathbf{1}_{n-c-1}\right)\mathsf{I}_{d}$ (57)
$=\tilde{\sigma}^{2}\mathsf{I}_{d}$ (58)
Proof:  [Proof of Claim 1] $\mathbf{Y}_{\ell}=\hat{M}(\mathbf{x}_{\ell})=\mathbf{x}_{\ell}+\sum_{j<\ell,j\notin\mathcal{C}}\mathbf{S}_{j,\ell}-\sum_{j>\ell,j\notin\mathcal{C}}\mathbf{S}_{\ell,j}+\mathbf{N}_{\ell}$, for all $\ell\notin\mathcal{C}$. Therefore, $\mathbf{Y}_{\ell}\sim\mathcal{N}(\mathbf{x}_{\ell},(\sigma^{2}+rc)\mathsf{I}_{d})$. The distribution of the $k$th component of $\mathbf{Y}_{\ell}$, across all $\ell\in\mathcal{C}$ is given by,
$\begin{pmatrix}Y_{1,k}\\ \vdots\\ Y_{n-c,k}\end{pmatrix}\sim\mathcal{N}\begin{pmatrix}\begin{pmatrix}x_{1,k}\\ \vdots\\ x_{n-c,k}\end{pmatrix}_{(n-c)\times 1},\begin{pmatrix}\sigma^{2}+rc&r&\dotsc&r\\ r&\sigma^{2}+rc&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots&\\ r&r&\dotsc&\sigma^{2}+rc\end{pmatrix}_{(n-c)\times(n-c)}\end{pmatrix}$ (59)
for $k\in[1:d]$, as $\text{cov}(Y_{i,k},Y_{j,k})=-\mathbb{E}[S_{i,j}^{2}]=r$ for $i\neq j$ and $\text{var}(Y_{i,k})=-r(n-c-1)+\sigma^{2}+r(n-1)=\sigma^{2}+rc$, $\forall i$. From the above distributions and from the fact that $\text{cov}(Y_{i,k},Y_{j,k})=0$, $\forall i,j\in[1:n]\setminus\mathcal{C}$ and $\forall k\neq k^{\prime}$, we derive,
$\begin{pmatrix}\mathbf{Y}_{1}\\ \vdots\\ \mathbf{Y}_{n-c}\end{pmatrix}\!\sim\!\mathcal{N}\begin{pmatrix}\!\!\begin{pmatrix}\mathbf{x}_{1}\\ \vdots\\ \mathbf{x}_{n-c}\end{pmatrix}_{\!\!\!(n\!-\!c)\times 1}\!\!\!,\begin{pmatrix}(\sigma^{2}+rc)\mathsf{I}_{d}&r\mathsf{I}_{d}&\dotsc&r\mathsf{I}_{d}\\ r\mathsf{I}_{d}&(\sigma^{2}+rc)\mathsf{I}_{d}&\dotsc&r\mathsf{I}_{d}\\ \vdots&\vdots&\vdots&\vdots&\\ r\mathsf{I}_{d}&r\mathsf{I}_{d}&\dotsc&(\sigma^{2}+rc)\mathsf{I}_{d}\end{pmatrix}_{\!\!\!(n\!-\!c)d\times(n\!-\!c)d}\end{pmatrix}$ (60)
We use the following standard result on the conditional distributions of multivariate Gaussian distributions [43]. Let $\mathbf{V}\in\mathbb{R}^{d}$ be $\mathbf{V}\sim\mathcal{N}(\mathbf{\hat{\mu}},\hat{\Sigma})$. Consider the partition $\mathbf{V}=[\mathbf{V}_{1},\mathbf{V}_{2}]^{T}$ with $\mathbf{V}_{1}\in\mathbb{R}^{p}$ and $\mathbf{V}_{2}\in\mathbb{R}^{d-p}$, and let the corresponding partitions of $\mathbf{\hat{\mu}}$ and $\hat{\Sigma}$ be $\mathbf{\hat{\mu}}=[\mathbf{\hat{\mu}}_{1},\mathbf{\hat{\mu}}_{2}]^{T}$ and $\hat{\Sigma}=\begin{pmatrix}\hat{\Sigma}_{1,1}&\hat{\Sigma}_{1,2}\\ \hat{\Sigma}_{2,1}&\hat{\Sigma}_{2,2}\end{pmatrix}$. Then, the conditional distribution of $\mathbf{V}_{1}|\mathbf{V}_{2}=\mathbf{v}_{2}$ is given by $\mathbf{V}_{1}|\mathbf{V}_{2}=\mathbf{v}_{2}\sim\mathcal{N}(\mathbf{\mu}_{*},\Sigma_{*})$ where,
$\mathbf{\mu}_{*}$ $=\mathbf{\hat{\mu}}_{1}+\hat{\Sigma}_{1,2}\hat{\Sigma}_{2,2}^{-1}(\mathbf{v}_{2}-\mathbf{\hat{\mu}}_{2})$ (61)
$\Sigma_{*}$ $=\hat{\Sigma}_{1,1}-\hat{\Sigma}_{1,2}\hat{\Sigma}_{2,2}^{-1}\hat{\Sigma}_{2,1}$ (62)
Based on this result, we have, $\mathbf{Y}_{1}|\mathbf{Y}_{j};j\in[2:n-c]\sim\mathcal{N}(\tilde{\mathbf{\mu}},\tilde{\Sigma})$ where,
$\tilde{\mathbf{\mu}}$ $=\mathbf{x}_{1}+(r\mathbf{1}_{n-c-1}^{T}\otimes\mathsf{I}_{d})\left(\begin{pmatrix}(\sigma^{2}+rc)&r&\dotsc&r\\ r&(\sigma^{2}+rc)&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots&\\ r&r&\dotsc&(\sigma^{2}+rc)\end{pmatrix}_{(n-c-1)\times(n-c-1)}\!\!\!\!\!\!\!\!\!\otimes\mathsf{I}_{d}\right)^{-1}$
×$\qquad\qquad\qquad\qquad\qquad\qquad\times\begin{pmatrix}\mathbf{y}_{2}-\mathbf{x}_{2}\\ \vdots\\ \mathbf{y}_{n-c}-\mathbf{x}_{n-c}\end{pmatrix}$ (63)
$=\mathbf{x}_{1}+\left(r\mathbf{1}^{T}_{n-c-1}\begin{pmatrix}(\sigma^{2}+rc)&r&\dotsc&r\\ r&(\sigma^{2}+rc)&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots&\\ r&r&\dotsc&(\sigma^{2}+rc)\end{pmatrix}_{(n-c-1)\times(n-c-1)}^{-1}\otimes\mathsf{I}_{d}\right)$
×$\qquad\qquad\qquad\qquad\qquad\qquad\times\begin{pmatrix}\mathbf{y}_{2}-\mathbf{x}_{2}\\ \vdots\\ \mathbf{y}_{n-c}-\mathbf{x}_{n-c}\end{pmatrix}$ (64)
using the properties $(A\otimes B)^{-1}=A^{-1}\otimes B^{-1}$ and $(A\otimes B)(C\otimes D)=(AC)\otimes(BD)$ of Kronecker products. Moreover,
$\tilde{\Sigma}$ $=(\sigma^{2}+rc)\mathsf{I}_{d}-(r\mathbf{1}_{n-c-1}^{T}\otimes\mathsf{I}_{d})\left(\!\!\!\begin{pmatrix}\sigma^{2}+rc&r&\dotsc&r\\ r&\sigma^{2}+rc&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots&\\ r&r&\dotsc&\sigma^{2}+rc\end{pmatrix}_{(n-c-1)\times(n-c-1)}\!\!\!\!\!\!\!\!\!\!\!\otimes\mathsf{I}_{d}\right)^{-1}$
×$\qquad\qquad\qquad\qquad\times(r\mathbf{1}_{n-c-1}\otimes\mathsf{I}_{d})$ (65)
$=(\sigma^{2}+rc)\mathsf{I}_{d}-r^{2}\mathbf{1}^{T}_{n-c-1}\begin{pmatrix}\sigma^{2}+rc&r&\dotsc&r\\ r&\sigma^{2}+rc&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots&\\ r&r&\dotsc&\sigma^{2}+rc\end{pmatrix}_{(n-c-1)\times(n-c-1)}^{-1}\!\!\!\!\!\!\!\!\!\!\!\!\mathbf{1}_{n-c-1}\mathsf{I}_{d}$ (66)
$=\left((\sigma^{2}+rc)-r^{2}\mathbf{1}^{T}_{n-c-1}\begin{pmatrix}\sigma^{2}+rc&r&\dotsc&r\\ r&\sigma^{2}+rc&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots&\\ r&r&\dotsc&\sigma^{2}+rc\end{pmatrix}_{(n-c-1)\times(n-c-1)}^{-1}\!\!\!\!\!\!\!\!\!\!\!\!\mathbf{1}_{n-c-1}\right)\mathsf{I}_{d}$ (67)
$=\tilde{\sigma}^{2}\mathsf{I}_{d}$ (68)
$\blacksquare$
Next, we apply Claim 1 in the privacy constraint in (54) to obtain the condition on $\sigma^{2}$ and $r$ to ensure $(\epsilon,\delta)$-DP.
Claim 2
The DP-DME system in Theorem 1 satisfies $(\epsilon,\delta)$-DP when,
$\tilde{\sigma}^{2}\geq\sigma^{2}_{\epsilon,\delta}.$ (69)
Proof:  [Proof of Claim 2] For given values of $\mathbf{Y}_{j}=\mathbf{y}_{j}$, $\forall j\in[2:n-c]$, consider the variation of $\mathbf{x}_{1}$ by fixing all $\mathbf{x}_{j}$, $\forall j\in[2:n-c]$, and define,
$f(\mathbf{x}_{1})$ $=\tilde{\mathbf{\mu}}$ (70)
$=\mathbf{x}_{1}+\left(r\mathbf{1}_{n-c-1}^{T}\begin{pmatrix}\sigma^{2}+rc&r&\dotsc&r\\ r&\sigma^{2}+rc&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots\\ r&r&\dotsc&\sigma^{2}+rc\end{pmatrix}_{n-c-1}^{-1}\otimes\mathsf{I}_{d}\right)\begin{pmatrix}\mathbf{y}_{2}-\mathbf{x}_{2}\\ \vdots\\ \mathbf{y}_{n-c}-\mathbf{x}_{n-c}\end{pmatrix}$ (71)
Then, define a new random variable $\mathbf{W}=f(\mathbf{x}_{1})+\mathbf{N}$, where $\mathbf{N}\sim\mathcal{N}\left(\mathbf{0}_{d},\tilde{\sigma}^{2}\mathsf{I}_{d}\right)$. Note that for any given values of $\mathbf{y}_{j}$, $\mathbf{x}_{j}$, $\forall j\in[2:n-c]$, $\mathbf{W}\sim\mathbf{Y}_{1}|\mathbf{Y}_{j}=\mathbf{y}_{j},j\in[2:n-c]$, (statistically equivalent). Now, consider the $(\epsilon,\delta)$-DP constraint in (54).
$\mathbb{P}(M(\mathbf{x}_{1})\in\mathcal{A}|$ $M(\mathbf{x}_{j})=\mathbf{y}_{j},j\in[2:n-c])\leq e^{\epsilon}\mathbb{P}(M(\mathbf{x}_{1}^{\prime})\in\mathcal{A}|M(\mathbf{x}_{j})=\mathbf{y}_{j},j\in[2:n-c])+\delta$ (72)
which is equivalent to,
$\mathbb{P}(f(\mathbf{x}_{1})+\mathbf{N}\in\mathcal{A})\leq e^{\epsilon}\mathbb{P}(f(\mathbf{x}^{\prime}_{1})+\mathbf{N}\in\mathcal{A})+\delta.$ (73)
As (73) represent the standard Gaussian mechanism in DP for the query $f(\mathbf{x}_{1})$, we use the results from [39, 40] to obtain the values of $\sigma^{2}$ and $r$ that satisfy (73). We restate Theorem 8 of [39] here for completeness.
Theorem 8 of [39] Let $f:\mathbb{X}\to\mathbb{R}^{d}$ be a function with global $L_{2}$ sensitivity $\Delta$. For any $\epsilon>0$ and $\delta\in[0,1]$ the Gaussian output perturbation mechanism $M(x)=f(x)+Z$ with $Z\sim\mathcal{N}(\mathbf{0}_{d},\hat{\sigma}^{2}\mathsf{I}_{d})$ is $(\epsilon,\delta)$-DP if and only if,
$\Phi\left(\frac{\Delta}{2\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{\Delta}\right)-e^{\epsilon}\Phi\left(-\frac{\Delta}{2\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{\Delta}\right)\leq\delta,$ (74)
where $\Phi(\cdot)$ is the CDF of the standard Gaussian distribution.
Applying Theorem 8 of [39] directly on (73) gives,
$\tilde{\sigma}^{2}\geq\sigma^{2}_{\epsilon,\delta},$ (75)
where $\sigma^{2}_{\epsilon,\delta}=\inf_{\hat{\sigma}>0}\{\hat{\sigma};\Phi\left(\frac{\Delta}{2\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{\Delta}\right)-e^{\epsilon}\Phi\left(-\frac{\Delta}{2\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{\Delta}\right)\leq\delta\}$ with $\Delta=2$. The value of $\Delta$ is calculated as,
$\Delta$ $=\sup_{\mathbf{x}_{1}\in\mathbb{B}^{d}}\|f(\mathbf{x}_{1})-f(\mathbf{x^{\prime}}_{1})\|=\|\mathbf{x}_{1}-\mathbf{x^{\prime}}_{1}\|=2$ (76)
The lower bound in (75) is due to the fact that $\Phi\left(\frac{\Delta}{2\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{\Delta}\right)-e^{\epsilon}\Phi\left(-\frac{\Delta}{2\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{\Delta}\right)$ for any fixed $\epsilon$ and $\Delta=2$ is a decreasing function in $\hat{\sigma}$, which is proved next.
Claim 3
$\Phi\left(\frac{1}{\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{2}\right)-e^{\epsilon}\Phi\left(-\frac{1}{\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{2}\right)$ is a decreasing function in $\hat{\sigma}$.
Proof:  Using the definition of the standard Gaussian CDF, for a fixed $\epsilon$ we have,
[…]
As $f(\hat{\sigma})$ is a smooth and continuous function, its derivative is given by,
$f^{\prime}(\hat{\sigma})$ $=\frac{1}{2\pi}\left(-\frac{1}{\sigma^{2}}-\frac{\epsilon}{2}\right)e^{-\frac{1}{2}\left(\frac{1}{\hat{\sigma}^{2}}+\frac{\epsilon^{2}\hat{\sigma}^{2}}{4}-\epsilon\right)}-\frac{e^{\epsilon}}{2\pi}\left(\frac{1}{\sigma^{2}}-\frac{\epsilon}{2}\right)e^{-\frac{1}{2}\left(\frac{1}{\hat{\sigma}^{2}}+\frac{\epsilon^{2}\hat{\sigma}^{2}}{4}+\epsilon\right)}$ (78)
$=-\frac{1}{\sigma^{2}\pi}e^{-\frac{1}{2}\left(\frac{1}{\hat{\sigma}^{2}}-\frac{\epsilon\hat{\sigma}}{2}\right)^{2}}<0.$ (79)
$\blacksquare$
This concludes the proof of Claim 2. $\blacksquare$
Substituting for $\tilde{\sigma}^{2}$ in (69) from (58) gives,
$\left(\sigma^{2}+rc-r^{2}\mathbf{1}_{n-c-1}^{T}\begin{pmatrix}\sigma^{2}+rc&r&\dotsc&r\\ r&\sigma^{2}+rc&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots\\ r&r&\dotsc&\sigma^{2}+rc\end{pmatrix}_{n-c-1}^{-1}\mathbf{1}_{n-c-1}\right)$ $\geq\sigma^{2}_{\epsilon,\delta}$ (80)
$\sigma^{2}+rc-r^{2}\mathbf{1}_{n-c-1}^{T}\big{(}(\sigma^{2}+r(c-1))\mathsf{I}_{n-c-1}+r\mathbf{1}_{n-c-1}\mathbf{1}_{n-c-1}^{T}\big{)}^{-1}\mathbf{1}_{n-c-1}$ $\geq\sigma^{2}_{\epsilon,\delta}$ (81)
Using the Sherman-Morrison formula, (81) simplifies to,
$\sigma^{2}+rc-r^{2}\mathbf{1}^{T}_{n-c-1}\left(\frac{1}{\sigma^{2}+r(c-1)}I_{n-c-1}-\frac{\frac{r}{(\sigma^{2}+r(c-1))^{2}}\mathbf{1}_{n-c-1}\mathbf{1}_{n-c-1}^{T}}{1+\frac{r}{\sigma^{2}+r(c-1)}\mathbf{1}_{n-c-1}^{T}\mathbf{1}_{n-c-1}}\right)\mathbf{1}_{n-1}$ $\geq\sigma^{2}_{\epsilon,\delta}$ (82)
$\frac{(\sigma^{2}+r(c-1))(\sigma^{2}+r(n-1))}{\sigma^{2}+r(n-2)}$ $\geq\sigma^{2}_{\epsilon,\delta}$ (83)
which is equivalent to the expression in (46). The analysis of the roots of (46) along with the constraint $\sigma^{2}+r(n-1)>0$ (for a positive variance of $\mathbf{N}_{i}$s) result in the values of $r$ given in (47), that makes the system satisfy $(\epsilon,\delta)$-DP for a fixed $\sigma^{2}$. It can be shown that for any fixed $\sigma^{2}<\sigma^{2}_{\epsilon,\delta}$, $\sigma^{2}+r(n-1)\leq 0$ for any $r$ satisfying (46). Matrices of the form $\Sigma=(\sigma^{2}+r(c-1))\mathsf{I}_{n}+r\mathbf{1}_{n}\mathbf{1}_{n}^{T}$ with such $\sigma$ and $r$ are not valid covariance matrices of $(Y_{1,k},\dotsc,Y_{n-c,k})$, $k\in[1:d]$, as they are not positive definite, i.e., $\mathbf{1}_{n-c}^{T}\Sigma\mathbf{1}_{n-c}=(n-c)(\sigma^{2}+r(n-1))\leq 0$. This verifies the requirement $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$.
Next, we prove that the covariance matrix of $[\mathbf{Y}_{1,k},\dotsc,\mathbf{Y}_{n-c,k}]^{T}$ for any $0\leq c<t$, $\forall k$, i.e., matrix $\Sigma=(\sigma^{2}+r(c-1))\mathsf{I}_{n-c}+r\mathbf{1}_{n-c}\mathbf{1}_{n-c}^{T}$ with any fixed $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$ and any corresponding $r$ in the range (47) is positive definite.
Claim 4
The matrix $\Sigma$ with $\Sigma_{i,i}=\sigma^{2}+rc$, $\forall i\in[1:n-c]$ and $\Sigma_{i,j}=r$, $\forall i,\in[1:n-c]$, $i\neq j$, is positive definite for any fixed $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$ and for any $r$ in the range (47).
Proof:  [Proof of Claim 4] From the properties of circulant matrices [44], the eigenvalues of $\Sigma$ are given by,
$\lambda_{j}$ $=\sigma^{2}+r\sum_{k=1}^{n-1}\omega^{kj},\quad j=0,\dotsc,n-1$ (84)
=$=\begin{cases}\sigma^{2}+r(n-1),&j=0\\ \sigma^{2}+r(c-1),&j=1,\dotsc,n-1\end{cases}$ (85)
where $\omega$ is the $n$th root of unity. It remains to prove that $\lambda_{j}>0$, $\forall j$ to prove the positive definiteness of $\Sigma$.
case 1: $c=0$: From direct substitution of the upper and lower bounds in (48) on $r$, one can observe that $\sigma^{2}-r>0$ and $\sigma^{2}+r(n-1)>0$, respectively. The explicit calculations are given below. This proves $\lambda_{j}>0$, $\forall j$.
Proof of $\sigma^{2}+r(n-1)>0$: Considering the lower bound on $r$ in (48), we have,
$\sigma^{2}+r(n-1)$
$\geq\sigma^{2}+\left(\frac{(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{2(n-1)}-\frac{\sqrt{(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}+4(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(n-1)}\right)(n-1)$ (86)
$=\frac{1}{2}\left(2\sigma^{2}+(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sqrt{(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}+4(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}\right)$ (87)
$=\frac{1}{2}(p-q),$ (88)
where $p=2\sigma^{2}+(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})>0$ and $q=\sqrt{(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}+4(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}\geq 0$. Note that,
$p^{2}-q^{2}$ $=4\sigma^{4}+4\sigma^{2}(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-4(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})$ (89)
$=4\sigma^{2}(\sigma^{2}-\sigma^{2}+\sigma^{2}_{\epsilon,\delta})>0,$ (90)
which implies that $p>q$.
Proof of $\sigma^{2}-r>0$: Considering the upper bound on $r$ in (48),
$\sigma^{2}-r$ $\geq\sigma^{2}-\frac{(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{2(n-1)}-\frac{\sqrt{(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}+4(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(n-1)}$ (91)
$=\tilde{p}-\tilde{q},$ (92)
where $\tilde{p}=\sigma^{2}-\frac{(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{2(n-1)}>0$ and $\tilde{q}=\frac{\sqrt{(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}+4(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(n-1)}\geq 0$. Note that $\tilde{p}^{2}-\tilde{q}^{2}>0$, which implies that $\tilde{p}>\tilde{q}$.
case 2: $c=1$: If $r\geq 0$, $\lambda_{j}>0$, $\forall j$, and $\Sigma$ is positive definite. For $r<0$ satisfying (47), $\sigma^{2}+r(c-1)\geq\sigma^{2}+r(n-1)$. Therefore, it remains to prove that $\sigma^{2}+r(n-1)>0$ for $r$ satisfying (47) for any $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$. Using the lower bound in (47),
$\sigma^{2}+r(n-1)\geq\sigma^{2}-\frac{\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})(n-1)}{\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2)}=\frac{\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2)}>0.$ (93)
case 3: $c>1$: Similar to case 2, if $r\geq 0$, $\lambda_{j}>0$, $\forall j$, and $\Sigma$ is positive definite. For $r<0$ satisfying (47), $\sigma^{2}+r(n-1)\leq\sigma^{2}+r(c-1)$. Therefore, it remains to prove that $\sigma^{2}+r(n-1)>0$ for $r$ satisfying (47) for any $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$. From the lower bound on $r$ in (47), we have,
$\sigma^{2}+r(n-1)$
$\geq\sigma^{2}+\frac{-(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c}{2(c-1)}+\frac{\sqrt{((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(c-1)}$ (94)
$\geq\frac{-((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)-2\sigma^{2}+\sqrt{((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(c-1)}$ (95)
If $-((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)-2\sigma^{2}\geq 0$, $\sigma^{2}+r(n-1)>0$ and the claim is proved. Consider the case where $-((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)-2\sigma^{2}<0$. To prove that $\sigma^{2}+r(n-1)>0$, we prove that $p>|q|$, where,
$p$ $=\sqrt{((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}$ (96)
$q$ $=-((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)-2\sigma^{2}$ (97)
As $p>0$ and $|q|>0$, consider,
$p^{2}-|q|^{2}$ $=((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})$
$\quad-((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}-4\sigma^{2}((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)-4\sigma^{4}$ (98)
$=-4\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})(c-1)+4\sigma^{4}(c-1)$ (99)
$=4\sigma^{2}\sigma^{2}_{\epsilon,\delta}(c-1)>0$ (100)
as $c>1$. This proves $p>|q|$. $\blacksquare$
Claims 1 and 2 collectively provide the feasible range of $r$ for a fixed $\sigma^{2}$ to ensure privacy and Claim 4 proves that these values of $\sigma^{2}$ and (corresponding) $r$ result in valid covariance matrices, which proves Lemma 1. $\blacksquare$
Next, we find the optimum values of $\sigma^{2}$ and $r$ that minimizes the MSE while satisfying the privacy constraint in Lemma 1.
Lemma 2 (Optimum noise parameters - No dropouts)
For any fixed $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$, the optimum $r$ that satisfies $(\epsilon,\delta)$-DP is given by,
$r_{*}$ $=\arg\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\sigma^{2},r,\bm{\alpha})$ (101)
=$=\begin{cases}a-b,&c=0\\ \frac{-\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2)},&c=1\\ \frac{-(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c}{2(n-1)(c-1)}+\frac{\sqrt{((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(n-1)(c-1)},&c>1\end{cases}$ (102)
where,
$a$ $=\frac{(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{2(n-1)}$ (103)
$b$ $=\frac{\sqrt{(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}+4(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(n-1)}.$ (104)
For the case where $t=n$, $\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\sigma^{2},r,\bm{\alpha})$ is a decreasing function in $\sigma^{2}$ and,
$\min_{\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}}\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\sigma^{2},r,\bm{\alpha})$ $=\lim_{\sigma^{2}\to\infty}\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\sigma^{2},r,\bm{\alpha})$ (105)
=$=\begin{cases}O\left(\frac{d\sigma^{2}_{\epsilon,\delta}}{(n-c)^{2}}\right),&\text{if }(n-c)^{2}\gg d\\ O(1),&otherwise.\end{cases}$ (106)
Proof:  [Proof of Lemma 2] The proof consists of two steps, namely, 1) optimize $r$ for fixed $\sigma^{2}$, 2) optimize $\sigma^{2}$ to minimize the MSE. The first step is characterized in Claim 5.
Claim 5
For any fixed $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$ the optimum $r$ is given in (102).
Proof:  For any given $\sigma^{2}$, $r$, $\mathcal{F}$ and $\mathcal{C}$, from Proposition 1,
$\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})$ $=\left(1+\frac{|\mathcal{U}|/d}{\sigma^{2}+r(|\mathcal{F}|-1)}\right)^{-1}$ (107)
is an increasing function in $r$. For a given $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$, any $r$ in the range specified in Lemma 1 satisfies $(\epsilon,\delta)$-DP, and results in a valid covariance matrix (Claim 4). Therefore, the optimum $r$ is given by the minimum value in the respective ranges (47)-(48). $\blacksquare$
To find the optimum $\sigma^{2}$, we begin with the following claim.
Claim 6
For the case of no dropouts, i.e., $t=n$, $\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{C},\sigma^{2},r,\bm{\alpha})$ is a decreasing function in $\sigma^{2}$ for any $\mathcal{C}$ satisfying $0\leq|\mathcal{C}|<n$.
Proof:  From Proposition 1 and Lemma 1,
$\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{C},\sigma^{2},r,\bm{\alpha})=\left(1+\frac{|\mathcal{U}|/d}{\sigma^{2}+r_{*}(n-1)}\right)^{-1}$ (108)
as $|\mathcal{F}|=n$. Let $f(\sigma^{2})=\sigma^{2}+r_{*}(n-1)$. We prove that (108) decreases with $\sigma^{2}$ by showing that $f(\sigma^{2})$ is a decreasing function in $\sigma^{2}$ for each of the three cases in (102).
case 1: $c=0$:
$\frac{df(\sigma^{2})}{d\sigma^{2}}$ $=1+(n-1)\frac{dr_{*}}{d\sigma^{2}}$ (109)
$=1+\frac{n-2}{2}-\frac{(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})(n-2)^{2}+2(n-1)(2\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{2\sqrt{(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}+4(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}$ (110)
$=\frac{1}{2}(\hat{p}-\hat{q})$ (111)
where $\hat{p}=n>0$ and $\hat{q}=\frac{(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})(n-2)^{2}+2(n-1)(2\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{\sqrt{(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}+4(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}>0$. As $\hat{p}^{2}-\hat{q}^{2}<0$ implies $\frac{df(\sigma^{2})}{d\sigma^{2}}<0$, consider:
$(\hat{p}^{2}-\hat{q}^{2})A^{2}$
$=n^{2}(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}+4n^{2}(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})$
$\quad-(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}(n-2)^{4}-4(n-1)^{2}(2\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}-4(n-1)(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})(2\sigma^{2}-\sigma^{2}_{\epsilon,\delta})$ (112)
$=4(n-1)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}(n-2)^{2}+4(n-1)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})(n^{2}\sigma^{2}-(n-2)^{2}(2\sigma^{2}-\sigma^{2}_{\epsilon,\delta}))$
$\quad-4(n-1)^{2}(2\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}$ (113)
$=4(n-1)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}(n-2)^{2}+16(n-1)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})\sigma^{2}$
$\quad-4(n-1)(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}-4(n-1)^{2}(2\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}$ (114)
$=4(n-1)^{2}(4\sigma^{4}-4\sigma^{2}\sigma^{2}_{\epsilon,\delta}-4\sigma^{4}+4\sigma^{2}\sigma^{2}_{\epsilon,\delta}-\sigma^{4}_{\epsilon,\delta})$ (115)
$=-4\sigma^{4}_{\epsilon,\delta}(n-1)^{2}<0$ (116)
where $A=\sqrt{(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}+4(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}$.
case 2: $c=1$:
$r_{*}$ $=\frac{-\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2)}$ (117)
$\frac{df(\sigma^{2})}{d\sigma^{2}}$ $=1+(n-1)\frac{dr_{*}}{d\sigma^{2}}$ (118)
$=1+(n-1)\left(\frac{(\sigma^{2}_{\epsilon,\delta}-2\sigma^{2})(\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2))+(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{(\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2))^{2}}\right)$ (119)
$=-\sigma^{4}_{\epsilon,\delta}(n-2)<0$ (120)
case 3: $c>1$:
$r*$ $=\frac{-(n\!-\!2)(\sigma^{2}\!-\!\sigma^{2}_{\epsilon,\delta})\!-\!\sigma^{2}c}{2(n\!-\!1)(c\!-\!1)}\!+\!\frac{\sqrt{((n\!-\!2)(\sigma^{2}\!-\!\sigma^{2}_{\epsilon,\delta})\!-\!\sigma^{2}c)^{2}+4(n\!-\!c\!-\!1)\sigma^{2}(\sigma^{2}\!-\!\sigma^{2}_{\epsilon,\delta})}}{2(n\!-\!1)(c\!-\!1)}$ (121)
$f(\sigma^{2})$ $=\sigma^{2}+r_{*}(n-1)$ (122)
$=\frac{\sigma^{2}(c\!-\!n)\!+\!\sigma^{2}_{\epsilon,\delta}(n\!-\!2)}{2(c\!-\!1)}\!+\!\frac{\sqrt{((n\!-\!2)(\sigma^{2}\!-\!\sigma^{2}_{\epsilon,\delta})-c\sigma^{2})^{2}\!+\!4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(c-1)}$ (123)
$\frac{df(\sigma^{2})}{d\sigma^{2}}$ $=\frac{c-n}{2(c-1)}+\frac{\sigma^{2}(n-c)^{2}-\sigma^{2}_{\epsilon,\delta}(n^{2}-nc-2n+2)}{2(c-1)\sqrt{((n\!-\!2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})\!-\!c\sigma^{2})^{2}+4(n\!-\!c\!\!-\!1)\sigma^{2}(\sigma^{2}\!-\!\!\sigma^{2}_{\epsilon,\delta})}}$ (124)
If $\sigma^{2}(n-c)^{2}-\sigma^{2}_{\epsilon,\delta}(n^{2}-nc-2n+2)\leq 0$, $\frac{df(\sigma^{2})}{d\sigma^{2}}<0$. Consider the case where $\sigma^{2}(n-c)^{2}-\sigma^{2}_{\epsilon,\delta}(n^{2}-nc-2n+2)>0$. Let $p=\frac{\sigma^{2}(n-c)^{2}-\sigma^{2}_{\epsilon,\delta}(n^{2}-nc-2n+2)}{2(c-1)\sqrt{((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-c\sigma^{2})^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}>0$ and $q=\frac{n-c}{2(c-1)}>0$, where $\frac{df(\sigma^{2})}{d\sigma^{2}}=p-q$. To prove that $\frac{df(\sigma^{2})}{d\sigma^{2}}<0$, it remains to prove $p^{2}-q^{2}<0$.
$p^{2}-q^{2}$ $=\frac{1}{4(c-1)^{2}}(B-(n-c)^{2})$ (125)
$=\frac{\sigma^{2}_{\epsilon,\delta}(n-1)}{B(c-1)}(c^{2}-nc+n-1)$ (126)
where $B=\frac{(\sigma^{2}(n-c)^{2}-\sigma^{2}_{\epsilon,\delta}(n^{2}-nc-2n+2))^{2}}{((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-c\sigma^{2})^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}>0$. Note that $c^{2}-nc+n-1<0$ for $22\leq c\leq n-$ 2. For $c=n-1$, $c^{2}-nc+n-1=0$. Therefore, $\frac{df(\sigma^{2})}{d\sigma^{2}}<0$ for $22\leq c\leq n-$ 2 and $\frac{df(\sigma^{2})}{d\sigma^{2}}=0$ when $c=n-1$, which corresponds to LDP. $\blacksquare$
Therefore, whenever $10\leq c<n-$ 1 $\min_{\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}}\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\sigma^{2},r,\bm{\alpha})$ is achieved when $\sigma^{2}\to\infty$. For $c=n-1$ which corresponds to the case of LDP, any $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$ (with the corresponding $r_{*}$) gives the same MSE. Hence, the solution $\sigma^{2}=\sigma^{2}_{\epsilon,\delta}$ and $r_{*}=0$ suffices to reach the MMSE when $c=n-1$.
Claim 7
For any $10\leq c<n-$ 1, $\lim_{\sigma^{2}\to\infty}\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{C},\sigma^{2},r,\bm{\alpha})=O\left(\frac{\sigma^{2}_{\epsilon,\delta}}{(n-c)^{2}}\right)$.
Proof:  Here we prove the claim for the most general case of $c>1$. The proof of other two cases $c=0$ and $c=1$ follow similar steps. Recall that for the case where $|\mathcal{F}|=n$,
$\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{C},\sigma^{2},r,\bm{\alpha})=\frac{1}{1+\frac{(n-c)/d}{\sigma^{2}+r_{*}(n-1)}}$ (127)
where,
$r_{*}=\frac{-(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c}{2(n-1)(c-1)}+\frac{\sqrt{((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(n-1)(c-1)}$ (128)
To find the limit of (127) as $\sigma^{2}\to\infty$, we first consider the following.
$\sqrt{((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}$
$\quad=(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{\frac{1}{2}}\left(\sigma^{2}(n-c)^{2}+\frac{c^{2}\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}-\sigma^{2}_{\epsilon,\delta}}-\sigma^{2}_{\epsilon,\delta}(n-2)^{2}\right)^{\frac{1}{2}}$ (129)
$\quad=\sigma\left(1-\frac{1}{2}\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}}-\frac{1}{4}\frac{1}{2!}\frac{\sigma^{4}_{\epsilon,\delta}}{\sigma^{4}}+\dotsc\right)\times\sigma(n-c)$
×$\quad\times\left(1+\frac{1}{2}\frac{\frac{c^{2}\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}-\sigma^{2}_{\epsilon,\delta}}-\sigma^{2}_{\epsilon,\delta}(n-2)^{2}}{\sigma^{2}(n-c)^{2}}-\frac{1}{4}\frac{1}{2!}\left(\frac{\frac{c^{2}\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}-\sigma^{2}_{\epsilon,\delta}}-\sigma^{2}_{\epsilon,\delta}(n-2)^{2}}{\sigma^{2}(n-c)^{2}}\right)^{2}+\dotsc\right)$ (130)
where (129) is obtained by applying the binomial expansion.
$\sigma^{2}+r_{*}(n-1)$
$=\frac{1}{2(c-1)}\left(\sigma^{2}(c-n)+\sigma^{2}_{\epsilon,\delta}(n-2)\right.$
$\qquad\qquad\qquad\left.+\sigma^{2}(n-c)\left(1-\frac{1}{2}\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}}+\frac{1}{2}\frac{\frac{c^{2}\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}-\sigma^{2}_{\epsilon,\delta}}-\sigma^{2}_{\epsilon,\delta}(n-2)^{2}}{\sigma^{2}(n-c)^{2}}+O(\frac{1}{\sigma^{4}})\right)\right)$ (131)
$=\frac{1}{2(c-1)}\left(\sigma^{2}_{\epsilon,\delta}(n-2)-\frac{1}{2}\sigma^{2}_{\epsilon,\delta}(n-c)+\frac{c^{2}\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{2(n-c)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}-\frac{\sigma^{2}_{\epsilon,\delta}(n-2)^{2}}{2(n-c)}+O(\frac{1}{\sigma^{2}})\right)$ (132)
Next, consider the limit,
$\lim_{\sigma^{2}\to\infty}\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{C},\sigma^{2},r,\bm{\alpha})$
$=\lim_{\sigma^{2}\to\infty}\left(1+\frac{(n-c)/d}{\sigma^{2}+r_{*}(n-1)}\right)^{-1}$ (133)
$=\lim_{\sigma^{2}\to\infty}\left(1\!+\!\frac{(n-c)/d}{\frac{1}{2(c-1)}\left(\sigma^{2}_{\epsilon,\delta}(n-2)\!-\!\frac{1}{2}\sigma^{2}_{\epsilon,\delta}(n-c)+\frac{c^{2}\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{2(n-c)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}-\frac{\sigma^{2}_{\epsilon,\delta}(n-2)^{2}}{2(n-c)}+O(\frac{1}{\sigma^{2}})\right)}\right)^{-1}$ (134)
$=\lim_{\sigma^{2}\to\infty}\left(1+\frac{(n-c)^{2}}{d\sigma^{2}_{\epsilon,\delta}}\right)^{-1}$ (135)
=$=\begin{cases}O\left(\frac{d\sigma^{2}_{\epsilon,\delta}}{(n-c)^{2}}\right),&(n-c)^{2}>>d\\ O(1),&otherwise.\end{cases}$ (136)
Claims 57 collectively prove Lemma 2. $\blacksquare$
$\blacksquare$
Next, we derive the optimum noise parameters for the case with dropouts, i.e., any $\mathcal{C}$ and $\mathcal{F}$ with $|\mathcal{F}|<n$. We first note that the DME process is the same as what is considered in the case with no dropouts with a reduced number of users. As the noise parameters of the privacy mechanism ($\sigma^{2}$ and $r$) are defined prior to the aggregation stage at the server, they are optimized for the worst case with the largest number of dropouts, resulting in an increase in the MSE compared to the case with no dropouts. However, the decoder can be optimized for the surviving number of users at the time of aggregation, as aggregation is performed after obtaining all responses from the remaining users. Therefore, the proof of the optimum noise parameters and the decoder consists of two steps, namely, 1) determining the optimum decoder for any given number of surviving users, 2) calculating the optimum noise parameters of the privacy mechanism, considering the worst case dropouts and colluding users.
Step 1 is direct from Proposition 1. Proposition 1 provides the optimum decoder for any fixed $\sigma^{2}$ and $r$ with any set of surviving users $\mathcal{F}\subset[1:n]$, $|\mathcal{F}|\geq t$. Even though all users may not respond when allowing for dropouts, the same conditions in Lemma 1 must be satisfied by the noise parameters to satisfy the strongest form of $(\epsilon,\delta)$-DP, considering the maximum information leakage that occurs when all $n$ users respond. Therefore, the problem to be solved is given by,
$\min_{r,\sigma^{2}}\!\!\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n\end{subarray}}\!\!\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})$ $=\min_{\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}}\!\!\min_{r}\!\!\!\!\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n\end{subarray}}\!\!\left(1+\frac{|\mathcal{U}|/d}{\sigma^{2}+r(|\mathcal{F}|-1)}\right)^{-1}$ (137)
$=\min_{\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}}\min_{r}\left(1+\frac{(t-c)/d}{\sigma^{2}+r(t-1)}\right)^{-1},$ (138)
as $\left(1+\frac{|\mathcal{U}|/d}{\sigma^{2}+r(|\mathcal{F}|-1)}\right)^{-1}$ is decreasing in $|\mathcal{F}|$ for any fixed $\sigma^{2}$ and $r$ that result in a valid covariance matrix, i.e., satisfies $\sigma^{2}+r(n-1)>0$. For any fixed $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$, $\left(1+\frac{(t-c)/d}{\sigma^{2}+r(t-1)}\right)^{-1}$ increases with $r$, and the optimal $r$ is given by the $r_{*}$ in (102). The next step is to solve,
$\sigma^{2}_{opt}$ $=\arg\min_{\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}}\min_{r}\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n\end{subarray}}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})$ (139)
$=\arg\min_{\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}}\left(1+\frac{(t-c)/d}{\sigma^{2}+r_{*}(t-1)}\right)^{-1}.$ (140)
Lemma 3 (Optimum noise variance - with dropouts)
For the case with $n$ users out of which up to any $n-t$ are allowed to dropout, the optimum noise variance $\sigma^{2}_{opt}$ that solves (139) is given by,
$\sigma^{2}_{opt}=\frac{\sigma^{2}_{\epsilon,\delta}(n^{2}-2n-cn+2)}{(n-c)^{2}}+\frac{\sigma^{2}_{\epsilon,\delta}(n-c-1)(n+c-2nc+t(n+c-2))}{(n-c)^{2}\sqrt{(t-c)(n-t)(n-c-1)}}.$ (141)
Proof:  To find $\sigma^{2}_{opt}$ consider the following notation for any given $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$ with $\epsilon,\delta,n,c$ and $t$ fixed.
$G(\sigma^{2})=\min_{r}\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n\end{subarray}}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})=\left(1+\frac{(t-c)/d}{\sigma^{2}+r_{*}(t-1)}\right)^{-1}$ (142)
Claim 8
The critical point of $G(\sigma^{2})$ is given by,
$\sigma^{2}_{cr}$ $=\frac{\sigma^{2}_{\epsilon,\delta}(n^{2}-2n-cn+2)}{(n-c)^{2}}+\frac{\sigma^{2}_{\epsilon,\delta}\lambda\sqrt{(n-c-1)(n-t)(t-c)}}{(n-c)^{2}(n-t)(t-c)}$ (143)
where $\lambda=(n+c)(t+1)-2(nc+t)$.
Proof: $G(\sigma^{2})=\left(1+\frac{(t-c)/d}{\sigma^{2}+r_{*}(t-1)}\right)^{-1}$ for any $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$ is a smooth and continuous function in $\sigma^{2}$ as $\sigma^{2}+r_{*}(t-1)$ is smooth and $\sigma^{2}+r_{*}(t-1)>0$ (proof in Section C, Claim 4). Consider the derivative of $G(\sigma^{2})$ with respect to $\sigma^{2}$ denoted by $G^{\prime}(\sigma^{2})$ to find its critical points.
$G^{\prime}(\sigma^{2})=-\left(\frac{-\frac{(t-c)}{d}\left(1+(t-1)\frac{dr_{*}}{d\sigma^{2}}\right)}{(\sigma^{2}+r_{*}(t-1))^{2}}\right)\left(1+\frac{(t-c)/d}{\sigma^{2}+r_{*}(t-1)}\right)^{-2}=0$ (144)
$0\displaystyle 1+(t-1)\frac{dr_{*}}{d\sigma^{2}}=$ 0 (145)
which simplifies to (143). $\blacksquare$
Next, we show that the critical point in Claim 8 is the global minimum of $G(\sigma^{2})$. Let $y=\frac{\sigma^{2}}{\sigma^{2}_{\epsilon,\delta}}$ and $y_{cr}=\frac{\sigma^{2}_{cr}}{\sigma^{2}_{\epsilon,\delta}}$.
case 1: $c>1$:
Recall that the $G^{\prime}(\sigma^{2})$ in (144) is of the form,
$G^{\prime}(\sigma^{2})$ $=K(t,c,\sigma^{2},r_{*})\left(1+(t-1)\frac{dr_{*}}{d\sigma^{2}}\right)$ (146)
$=K(t,c,\sigma^{2},r_{*})\left(1+\frac{t-1}{2(n-1)(c-1)}\left(-(n+c-2)+\frac{1}{2}\frac{1}{\sqrt{A(y)}}\frac{dA(y)}{dy}\right)\right)$ (147)
where $K(t,c,\sigma^{2},r_{*})=\frac{(t-c)/d}{(\sigma^{2}+r_{*}(t-1))^{2}}\left(1+\frac{(t-c)/d}{\sigma^{2}+r_{*}(t-1)}\right)^{-2}>0$ and
$A(y)=((n-2)(y-1)-yc)^{2}+4(n-c-1)y(y-1)$ (148)
Let $\phi(y)=\sqrt{A(y)}$, and Define,
$g(y)$ $=1+\frac{t-1}{2(n-1)(c-1)}\left(-(n+c-2)+h(y)\right)$ (149)
$h(y)$ $=\frac{1}{2}\frac{1}{\sqrt{A(y)}}\frac{dA(y)}{dy}$ (150)
$=\frac{(y(n-c-2)-(n-2))(n-c-2)-2(-n+c+1)(2y-1)}{\phi(y)}$ (151)
where $g(y_{cr})=0$. Note that $h(y)=\phi^{\prime}(y)$. Consider the derivative of $h(y)$.
$h^{\prime}(y)$ $=\frac{((n-c-2)^{2}+4(n-c-1))\phi(y)-\phi^{\prime}(y)h(y)\phi(y)}{\phi^{2}(y)}$ (152)
$=\frac{(n-c-2)^{2}+4(n-c-1)-h^{2}(y)}{\phi(y)}$ (153)
$=\frac{4(c-1)(n^{2}-n(c+2)+(c+1))}{\phi(y)}$ (154)
$=\frac{4(c-1)(n-c-1)(n-1)}{\phi(y)}\geq 0$ (155)
where the equality in (155) holds when $c=n-1$, which corresponds to the case of LDP for which $r_{*}=0$ and $\sigma^{2}_{*}=\sigma^{2}_{\epsilon,\delta}$. Thus, for $11<c<n-$ 1, $h^{\prime}(y)>0$. Then, for any arbitrarily small $\tilde{\delta}>0$,
$G^{\prime}(\sigma^{2}_{cr}+\tilde{\delta})$ $=K(t,c,\sigma^{2}_{cr}+\tilde{\delta},r_{*})\left(1+\frac{t-1}{2(n-1)(c-1)}\left(-(n-c-2)+h(\frac{\sigma^{2}_{cr}+\tilde{\delta}}{\sigma^{2}_{\epsilon,\delta}})\right)\right)$ (156)
$=K(t,c,\sigma^{2}_{cr}+\tilde{\delta},r_{*})\left(1+\frac{t-1}{2(n-1)(c-1)}\left(-(n-c-2)+h(y_{cr})+\delta\right)\right)$ (157)
$=K(t,c,\sigma^{2}_{cr}+\tilde{\delta},r_{*})\frac{\delta(t-1)}{2(n-1)(c-1)}>0$ (158)
Similarly,
$G^{\prime}(\sigma^{2}_{cr}-\tilde{\delta})$ $=K(t,c,\sigma^{2}_{cr}-\tilde{\delta},r_{*})\left(1+\frac{t-1}{2(n-1)(c-1)}\left(-(n-c-2)+h(\frac{\sigma^{2}_{cr}-\tilde{\delta}}{\sigma^{2}_{\epsilon,\delta}})\right)\right)$ (159)
$=K(t,c,\sigma^{2}_{cr}-\tilde{\delta},r_{*})\left(1+\frac{t-1}{2(n-1)(c-1)}\left(-(n-c-2)+h(y_{cr})-\delta\right)\right)$ (160)
$=K(t,c,\sigma^{2}_{cr}-\tilde{\delta},r_{*})\frac{-\delta(t-1)}{2(n-1)(c-1)}<0$ (161)
This proves that $\sigma^{2}_{cr}$ is a minimum of $G(\sigma^{2})$.
case 2: $c=1$:
When $c=1$, $y_{cr}$ corresponding to (8) is given by,
$y_{cr}=\frac{n-2}{n-1}+\frac{1}{n-1}\sqrt{\frac{(n-2)(t-1)}{n-t}}$ (162)
Recall that the $G^{\prime}(\sigma^{2})$ in (144) is of the form,
$G^{\prime}(\sigma^{2})$ $=K(t,c,\sigma^{2},r_{*})\left(1+(t-1)\frac{dr_{*}}{d\sigma^{2}}\right)$ (163)
$=K(t,c,\sigma^{2},r_{*})\left(1-\frac{t-1}{n-1}\left(1+\frac{n-2}{(y(n-1)-(n-2))^{2}}\right)\right)$ (164)
Let $\tilde{g}(y)=1-\frac{t-1}{n-1}(1+\tilde{h}(y))$ and $\tilde{h}(y)=\frac{n-2}{(y(n-1)-(n-2))^{2}}$. Note that, $h^{\prime}(y)=-\frac{2(n-1)(n-2)}{(y(n-1)-(n-2))^{3}}$, and $h^{\prime}(y)<0$ for $y>y_{cr}-\frac{1}{n-1}\sqrt{\frac{(n-2)(t-1)}{n-t}}$. Therefore, for any small $\delta\in(0,\frac{\sigma^{2}_{\epsilon,\delta}}{n-1}\sqrt{\frac{(n-2)(t-1)}{n-t}})$,
$G^{\prime}(\sigma^{2}_{cr}+\delta)$ $=K(t,c,\sigma^{2},r_{*})\left(1-\frac{t-1}{n-1}\left(1+h(\frac{\sigma^{2}_{cr}+\delta}{\sigma^{2}_{\epsilon,\delta}})\right)\right)$ (165)
$=K(t,c,\sigma^{2},r_{*})\left(1-\frac{t-1}{n-1}\left(1+h(y_{cr})-\tilde{\delta}\right)\right)$ (166)
$=K(t,c,\sigma^{2},r_{*})\left(\frac{\tilde{\delta}(t-1)}{n-1}\right)>0$ (167)
Similarly,
$G^{\prime}(\sigma^{2}_{cr}-\delta)$ $=K(t,c,\sigma^{2},r_{*})\left(1-\frac{t-1}{n-1}\left(1+h(\frac{\sigma^{2}_{cr}-\delta}{\sigma^{2}_{\epsilon,\delta}})\right)\right)$ (168)
$=K(t,c,\sigma^{2},r_{*})\left(1-\frac{t-1}{n-1}\left(1+h(y_{cr})+\tilde{\delta}\right)\right)$ (169)
$=K(t,c,\sigma^{2},r_{*})\left(\frac{-\tilde{\delta}(t-1)}{n-1}\right)<0,$ (170)
which proves that $\sigma^{2}_{cr}$ is a minimum of $G(\sigma^{2})$.
case 3: $c=0$:
As the $r_{*}$ is the same for both $c>1$ and $c=0$, proving that $\sigma^{2}_{cr}$ is the minimum of $G(\sigma^{2})$ follows the same steps as case 1. $\blacksquare$
As there are no other critical points satisfying $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$, the optimum $\sigma^{2}$ is given by $\sigma^{2}_{cr}$ and is presented in its complete form in (141). $\blacksquare$
Appendix D Proof of Proposition 2
Proposition 2 restated:
The following upper bounds hold for $\sigma^{2}_{\epsilon,\delta}$.
$\sigma^{2}_{\epsilon,\delta}$$\leq\begin{cases}\frac{8\ln(1.25/\delta)}{\epsilon^{2}},&\epsilon,\delta\in(0,1)\\ \frac{2\eta^{2}}{\epsilon},&\epsilon\geq 1,\delta\in(0,1)\end{cases}$ (171)
where
$\eta=\begin{cases}1+2\sqrt{\ln\frac{1}{2\delta}},&0<\delta\leq 0.05\\ 1+2\sqrt{\ln 10},&0.05<\delta<1.\end{cases}$ (172)
Proof:  We use the upper bounds on $\sigma_{\epsilon,\delta}$ derived in [40] and [8].
$\sigma_{\epsilon,\delta}$$\leq\begin{cases}\frac{2}{\epsilon}\sqrt{2\ln\frac{1.25}{\delta}},&\epsilon,\delta\in(0,1)\\ \frac{2}{\epsilon}\sqrt{2\ln\frac{1}{2\delta}}+\sqrt{\frac{2}{\epsilon}},&\epsilon\geq 1,\delta\in(0,0.05].\end{cases}$ (173)
Note that in the second case of (173), $\frac{2}{\epsilon}\sqrt{2\ln\frac{1}{2\delta}}+\sqrt{\frac{2}{\epsilon}}\leq\sqrt{\frac{2}{\epsilon}}\left(2\sqrt{\ln\frac{1}{2\delta}}+1\right)$ as $\epsilon\geq 1$. Moreover, it is clear that any $\hat{\sigma}$ satisfying $\Phi\left(\frac{1}{\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{2}\right)-e^{\epsilon}\Phi\left(-\frac{1}{\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{2}\right)\leq\delta_{1}$ also satisfies $\Phi\left(\frac{1}{\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{2}\right)-e^{\epsilon}\Phi\left(-\frac{1}{\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{2}\right)\leq\delta_{2}$, whenever $\delta_{1}\leq\delta_{2}$. Therefore, for any $\epsilon\geq 1$ and $\delta\in(0.05,1]$ we have,
$\sigma_{\epsilon,\delta}\leq\sigma_{\epsilon,0.05}\leq\sqrt{\frac{2}{\epsilon}}\left(2\sqrt{\ln 10}+1\right),$ (174)
where the first inequality comes from the definition of $\sigma_{\epsilon,\delta}$, given by $\sigma_{\epsilon,\delta}=\inf_{\hat{\sigma}>0}\{\hat{\sigma};\Phi\left(\frac{1}{\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{2}\right)-e^{\epsilon}\Phi\left(-\frac{1}{\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{2}\right)\leq\delta\}$. This determines the values of $\eta$ in (172).
$\blacksquare$
Appendix E Proof of Proposition 3
The proof of Proposition 3 is direct from the proof of Theorem 1, as in Theorem 1, we essentially solved
$(\sigma^{2}_{*},\rho_{*})=\arg\min_{\sigma^{2},\rho}\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n\end{subarray}}\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\sigma^{2}(1+\rho(|\mathcal{F}|-1),$ (175)
with $|\mathcal{C}|=c$, and then applied the decoder in Proposition 1. Note that (175) is equivalent to (17), which shows that the optimum noise parameters do not change. However, the resulting MSE changes as the decoder further normalizes the direct MSE resulted by the parameters in (175).
Appendix F Additional Details on Section 3
In this section, we go over the rigorous proofs of the geometric interpretation provided for the two-user example in Section 3.
Privacy Constraint:
For this example, (3) simplifies to,
$\mathbb{P}(\mathbf{x}_{i}+\mathbf{Z}_{i}\in\mathcal{A}|\mathbf{x}_{j}+\mathbf{Z}_{j}=\mathbf{y}_{j})\leq e^{\epsilon}\mathbb{P}(\mathbf{x}_{i}^{\prime}+\mathbf{Z}_{i}\in\mathcal{A}|\mathbf{x}_{j}+\mathbf{Z}_{j}=\mathbf{y}_{j})+\delta$ (176)
for each $i\neq j$, for any fixed $\mathbf{x}_{j},\mathbf{y}_{j}$ and $\forall\mathbf{x}_{i},\mathbf{x}_{i}^{\prime}$. Let $\mathbf{Y}_{i}=\mathbf{x}_{i}+\mathbf{Z}_{i}$ for $i=1,2$. The first step is to quantify the conditional distribution $\mathbf{Y}_{i}|\mathbf{Y}_{j}=\mathbf{y}_{j}$ for $i\neq j$. WLOG assume that $i=1$.
$\begin{pmatrix}\mathbf{Y}_{1}\\ \mathbf{Y}_{2}\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}\mathbf{x}_{1}\\ \mathbf{x}_{2}\end{pmatrix},\begin{pmatrix}\sigma^{2}\mathsf{I}_{d}&\rho\sigma^{2}\mathsf{I}_{d}\\ \rho\sigma^{2}\mathsf{I}_{d}&\sigma^{2}\mathsf{I}_{d}\end{pmatrix}\right)$ (177)
Then, from the theorems of conditional Gaussian distributions, we have, $\mathbf{Y}_{1}|\mathbf{Y}_{2}=\mathbf{y}_{2}\sim\mathcal{N}(\mathbf{\mu}_{*},\Sigma_{*})$ where,
$\mathbf{\mu}_{*}$ $=\mathbf{x}_{1}+\rho\mathsf{I}_{d}(\mathbf{y}-\mathbf{x}_{2})$ (178)
$\Sigma_{*}$ $=\sigma^{2}\mathsf{I}_{d}-\rho^{2}\sigma^{2}\mathsf{I}_{d}=\sigma^{2}(1-\rho^{2})\mathsf{I}_{d}=\tilde{\sigma}^{2}\mathbf{I}_{d}$ (179)
Define $f(\mathbf{x}_{i})=\mathbf{x}_{1}+\rho\mathsf{I}_{d}(\mathbf{y}-\mathbf{x}_{2})$. Then, define a new random variable $\mathbf{W}=f(\mathbf{x}_{1})+\mathbf{N}$, where $\mathbf{N}\sim\mathcal{N}\left(\mathbf{0}_{d},\tilde{\sigma}^{2}\mathsf{I}_{d}\right)$. Note that for any given values of $\mathbf{y}_{2}$, $\mathbf{x}_{2}$, $\mathbf{W}\sim\mathbf{Y}_{1}|\mathbf{Y}_{2}=\mathbf{y}_{2}$, (statistically equivalent). Now, consider the $(\epsilon,\delta)$-DP constraint in (176), which is equivalent to,
$\mathbb{P}(f(\mathbf{x}_{1})+\mathbf{N}\in\mathcal{A})\leq e^{\epsilon}\mathbb{P}(f(\mathbf{x}^{\prime}_{1})+\mathbf{N}\in\mathcal{A})+\delta.$ (180)
As (180) represents the standard Gaussian mechanism on $f(\mathbf{x}_{i})$ (with a sensitivity of $\sup_{\mathbf{x},\mathbf{x}^{\prime}}\|f(\mathbf{x}_{i})-f(\mathbf{x}_{i}^{\prime})\|_{2}=\|\mathbf{x}_{i}-\mathbf{x}_{i}^{\prime}\|=2$), the variance of $N$ must satisfy,
$\tilde{\sigma}^{2}\geq\sigma^{2}_{\epsilon,\delta}$ (181)
to ensure $(\epsilon,\delta)$-DP, where $\sigma^{2}_{\epsilon,\delta}=\inf_{\hat{\sigma}>0}\{\hat{\sigma};\Phi\left(\frac{\Delta}{2\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{\Delta}\right)-e^{\epsilon}\Phi\left(-\frac{\Delta}{2\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{\Delta}\right)\leq\delta\}$ with $\Delta=2$. This simplifies to,
$\sigma^{2}(1-\rho^{2})$ $\geq\sigma^{2}_{\epsilon,\delta}$ (182)
$\implies\quad\sigma\sin\theta$ $\geq\sigma_{\epsilon,\delta}$ (183)
Recall that $\mathbb{E}[\|\mathbf{Z}_{i}\|^{2}]=\sigma^{2}d=|\mathbf{Z}_{i}|^{2}$ from the vector representation explained in Section 3. Therefore, the privacy constraint in (176) simplifies to,
$|\mathbf{Z}_{i}|\sin\theta\geq\sigma_{\epsilon,\delta}\sqrt{d}=\hat{\sigma}_{\epsilon,\delta}$ (184)
which can be interpreted as the component of $\mathbf{Z}_{i}$ that is orthogonal to $\mathbf{Z}_{j}$ for $i\neq j$ having a variance that is lower bounded by a constant to ensure $(\epsilon,\delta)$-DP.
Optimum noise parameters:
The proof of $\sigma^{2}_{*}\to\infty$ and $\rho_{*}\to-1$ is given in the general proof of Lemma 2 in Appendix C, with $c=0$.